### v1

# Fast and Accurate Collimator-Detector Response Compensation in High-Energy SPECT Imaging with 1D Convolutions and Rotations

Abstract
Objective: Modeling of the collimator-detector response (CDR) in SPECT reconstruction enables improved resolution and more accurate quantitation, especially for higher energy imaging (e.g. 177Lu and 225Ac). Such modeling, however, can pose a significant computational bottleneck when there are substantial components of septal penetration and scatter in the acquired data, since a direct convolution-based approach requires large 2D kernels. The present work presents an alternative method for fast and accurate CDR compensation using a linear operator built from 1D convolutions and rotations (1D-R). To enable open-source development and use of these models in image reconstruction, we release a SPECTPSFToolbox repository for the PyTomography project on GitHub. Approach: A 1D-R CDR model was formulated, and subsequently fit to Monte Carlo 440 keV point source data representative of emissions in 225Ac imaging. Computation times of (i) the proposed 1D-R model and (ii) a traditional model that uses 2 dimensional convolutions (2D) were compared for typical SPECT matrix sizes. Both CDR modeling techniques were then used to reconstruct 225Ac phantom and patient data, and were compared by quantifying total counts in hot regions of interest (ROIs) and activity contrast between hot ROIs and background regions. Results: The 1D-R and 2D CDR models were created using the SPECTPSFToolbox. For typical matrix sizes in SPECT reconstruction, application of the 1D-R model provides a two-fold computational speed-up over the 2D model. Only small differences between the 1D-R and 2D models (order of $1\%$ ) were obtained for count and contrast quantification in select ROIs. Significance: A technique for CDR modeling in SPECT was proposed that (i) significantly speeds up reconstruction times, and (ii) yields nearly identical reconstructions to traditional 2D convolution based CDR techniques. The released toolbox will permit open-source development of similar models for different isotopes and collimators.
1 Introduction
Single photon emission computed tomography (SPECT) imaging is an in vivo modality of significant value in different clinical applications [1, 2, 3]. In particular, quantitative SPECT imaging holds great potential and value in the field of theranostics, wherein image quantification and dosimetry can be performed to improve radiopharmaceutical therapies (RPTs). Routine SPECT-based dosimetry could unveil the relationships between tumor-response and healthy organ complications with absorbed dose, therefore enabling personalized treatments that maximize tumor dose while minimizing toxicity to organs at risk [4, 5]. As a recent and well-known example, the success of 177Lu-PSMA-617 based RPTs in the VISION [6] and TheraP [7] randomized control trials has lead to expanded research and clinical translation efforts for wide-scale deployment, as well as pursuit of absorbed dose estimation as a tool towards personalized treatments [8, 9].
SPECT-based dosimetry relies on accurate quantification of activity distribution from SPECT images, which is strongly dependent on the system model used in image reconstruction. The system model is a linear operator that predicts the expectation of the acquired data given a 3D isotope distribution, accounting for phenomena such as photon attenuation in the patient. An important aspect of the model is collimator detector response (CDR) modeling, which estimates image blurring caused by the collimator and the scintillation crystals [10, 11, 12]. Computation of the CDR is typically the most computationally expensive operation in image reconstruction.
The CDR can be characterized by the detector response from a point source of activity; this is denoted the point spread function (PSF). The PSF can be decomposed into multiple components. The intrinsic response function (IRF) characterizes the inability of the scintillation crystals to precisely localize the point of interaction, and is sufficiently modeled using a Gaussian function. The collimator response for parallel hole collimators results from the inability of the collimator to accept only photons travelling perpendicular to the detector; it consists of three components: (i) the geometric response function (GRF) [11, 13, 14, 15], which describes photons that travel through the collimator holes without penetrating or interacting with the septa, (ii) the septal penetration response function (SPRF), which describes the contribution from photons that travel through the collimator without being attenuated, and (iii) the septal scatter response function (SSRF), which consists of photons that interacted and scattered within the collimator and were subsequently detected in the scintillator.
Selection of collimator parameters is an important aspect of SPECT imaging [16]. As the collimator becomes thicker and the diameter of the collimator bores becomes narrower, the relative contribution from the SPRF and SSRF decreases due to the increased attenuation probability for photons not travelling perpendicular to the detector. In this case, (i) the detector resolution improves and (ii) the net point spread function (PSF) is dominated by the IRF + GRF and can be reasonably approximated using a 2D Gaussian function. A trade-off of having a collimator with high septal thickness and small hole diameters, however, is a decrease in detector sensitivity; the corresponding implications for quantitative imaging is a decrease in precision or longer patient scan times. This trade-off must be independently considered for each isotope. For 177Lu labeled radiopharmaceuticals, where the imaged photons are 208208208~{}208keV, a commercially labeled “medium energy” collimator configuration (i) yields a reasonable count rate and (ii) adequately minimizes the SPRF and SSRF components. In this situation, a 2D Gaussian PSF model is sufficient for image reconstruction [17, 18].
Many commercially available reconstruction software only offer Gaussian PSF modeling, and thus implicitly assume the SPRF and SSRF are negligible. There are multiple advantages to this. Firstly, the computational advantage of 2D Gaussian PSF modeling is that 2D Gaussian convolution is separable into use of two perpendicular 1D Gaussian convolutions, which are less computationally expensive to implement. Secondly, the Gaussian PSF model used to model the GRF+IRF can be obtained analytically for any photon emission energy and standard collimator shapes, and thus does not require lookup tables. However, when the PSF has significant contributions from the SPRF and SSRF, 2D Gaussian PSF modeling fails to capture all the features of the PSF. This is typically an issue with radioisotopes that emit high photon energies, such as $\alpha$-emitters like 225Ac.
Use of $\alpha$-emitters in radiopharmaceutical therapies presents a major and exciting frontier, due to the high linear energy transfer (LET) associated with $\alpha$ emissions [19, 20, 21, 22, 23, 24]. In preclinical studies, simultaneous treatment with 177Lu-PSMA-617 and 225Ac-PSMA-617 compared to 177Lu-PSMA-617 alone resulted in significantly reduced tumor growth [25]. Kratochwil et al. [26] applied 225Ac-PSMA-617 to patients with metastatic castration resistant prostate cancer (mCRPC) who previously exhausted 177Lu-PSMA-617 treatment. Certain patients achieved a full response, with prostate specific antigen (PSA) levels in one patient decreasing from 419 ng/mL to below 0.1 ng/mL. At the time of writing, there are ongoing studies of 225Ac based radiopharmaceuticals looking at dose escalation [27, 28], fractionation [29], and safety and efficacy [30]. A recent meta analysis [31] elaborates further on these clinical trials, discussing toxicities and other challenges with these treatments. Quite recently, targeted $\alpha$ therapy with 213Bi has been shown to reduce amyloid plaque concentration in male mice [32], eluding to a potential treatment option for Alzheimer disease.
Throughout the decay chain of 225Ac, the daughters 213Bi and 221Fr emit photons detectable within a SPECT system of 440440440~{}440keV and 218218218~{}218keV respectively. Unfortunately, even with the commercially available collimators designed for high-energy photons, there are still significant SPRF and SSRF components present in the 213Bi 440440440~{}440keV peak; a sample PSF is shown in Figure 1. This phenomena similarly occurs in imaging of photons with energies of 511 keV (positron emitter) [33] and $364$ keV (131I) [34]. As a consequence, the PSF can no longer be modeled using a 2D Gaussian and is thus no longer separable into two perpendicular 1D components. Since this is the bottleneck of system modeling, image reconstruction takes significantly longer.
The reduction of reconstruction times in medical imaging remains an important research topic. Tsai et al. [35] recently showed that the limited-memory Broyden-Feltcher-Goldfarb-Shannon algorithm with box constraints and a diagonal preconditioner (L-BFGS-B-PC) was able to converge several times faster than the one step late expectation maximum (OSL-EM) algorithm in Positron Emission Tomography (PET) reconstruction because less projection operations were required. As another example, in Chun et al. [34], long reconstruction times for PSF modeling of 131I were partially remedied by using fast fourier transform (FFT) based 2D convolutions. By contrast, in the present work, we seek to reduce reconstruction times by using a PSF model that incorporates (i) 1D convolutions and (ii) rotations. It will be shown that this model significantly reduces computational time compared to standard and FFT based 2D convolutions, and consequently significantly reduces the time required for SPECT image reconstruction of 225Ac. The method is evaluated on phantom data and patient data, and we show that the proposed model yields identical results to Monte Carlo based 2D convolution techniques. All models considered in this paper are implemented using the GPU-accelerated functionality of PyTorch.
Alongside this paper, we release the SPECTPSFToolbox: an open-source GitHub repository which forms a new component of the PyTomography [36] project. The toolbox contains functionality for developing and fitting arbitrary PSF models to arbitrary point-source data. The saved models can then be loaded in our in-house initiated and community developed library PyTomography for SPECT reconstruction. As such, the techniques used in this paper can also be applied to other isotope / collimator configurations that are of interest in the nuclear medicine community. To encourage community use, we have released nine tutorials demonstrating how to use the toolbox, and how to integrate the models in PyTomography. The link to the PyTomography project (which includes the SPECTPSFToolbox) is https://github.com/PyTomography
2 Materials and Methods
In Section 2.1, the mathematical formalism of the PSF model is outlined, and the reconstruction protocol for all acquired data in subsequent sections is established. Section 2.2 describes how the model is fit to SIMIND [37] Monte Carlo 225Ac point source data acquisitions simulated at various source-detector distances. The computational time of the model is benchmarked and compared to 2D PSF modeling and Gaussian PSF modeling. In Section 2.3, the model is then used for reconstruction of (i) an 225Ac phantom consisting of spherical inserts and (ii) a patient receiving 225Ac-PSMA-617 treatment for metastatic prostate cancer. All computation was performed using a Microsoft Azure virtual machine (Standard NC6s v3) with a 6 CPUs (Intel(R) Xeon(R) CPU E5-2690 v4 @ 2.60GHz), 112 GB of RAM, and a Tesla V100 GPU.
2.1 Theory
2.1.1 PSF Modeling
In this work, the notation $f(x,y;d)$ is used for 3D objects: $d$ denotes the distance between a plane parallel to the detector and the detector, and $(x,y)$ denote the position on the plane. It is assumed that $x$, $y$, and $d$ are discrete and thus $f$ consists of voxels. The following notation is used to represent convolution operator $K$:
$K^{(x;d;b)}f\equiv\sum_{x^{\prime}}f(x^{\prime},y;d)k(x-x^{\prime};d;b)$ (1)
$K^{(y;d;b)}f\equiv\sum_{y^{\prime}}f(x,y^{\prime};d)k(y-y^{\prime};d)$ (2)
$K^{(x,y;d;b)}f\equiv\sum_{x^{\prime},y^{\prime}}f(x^{\prime},y^{\prime};d;b)k(x-x^{\prime},y-y^{\prime};d;b)$ (3)
where $b$ are additional parameters the kernel $k$ depends on, such as collimator septal thickness $L_{b}$, hole diameter $w_{b}$, and the linear attenuation coefficient of the collimator material $\mu_{b}$.
Assuming a linear shift invariant (LSI) PSF, the SPECT system matrix estimates the projection $g_{\phi}$ at angle $\phi$ as
$g_{\phi}(x,y)=\sum_{d}\sum_{x^{\prime},y^{\prime}}k_{\mathrm{PSF}}(x-x^{\prime},y-y^{\prime},d)p_{\mathrm{att}}(x,y,x^{\prime},y^{\prime},d)f(x,y,d)$ (4)
where $(x^{\prime},y^{\prime})$ is the source position on a plane parallel to the detector in 3D space, $\mathrm{psf}(...,d)$ is a kernel that yields the point spread function (PSF) at a distance $d$ from the detector, and $p_{\mathrm{att}}(x,y,x^{\prime},y^{\prime},d)$ is the probability that photons traveling from $(x,y,d)$ to detector coordinate $(x^{\prime},y^{\prime})$ are not attenuated. Under the assumption that the attenuation probabilities vary little across the PSF, it follows that $p_{\mathrm{att}}(x,y,x^{\prime},y^{\prime},d)\approx p_{\mathrm{att}}(x,y,x,y,d)$ and it can simply be expressed as $p_{\mathrm{att}}(x,y,d)$. Defining $f^{\prime}\equiv p_{\mathrm{att}}f$ as the “attenuation-adjusted” image, Equation 4 can then be rewritten in operator form (Equation 3) as
$g_{\phi}(x,y)=\sum_{d}K_{\mathrm{PSF}}^{(x,y;d;b)}f^{\prime}$ (5)
Since this convolution operation is typically the bottleneck of SPECT system matrix modeling and image reconstruction, it is of interest to look for techniques to reduce the computation time. Under conditions of no septal penetration and scatter, the CDR is dominated by the GRF: $K_{\mathrm{PSF}}^{(x,y;d;b)}$ can then be sufficiently approximated using a 2D Gaussian convolution $K_{G}^{(x,y;d;b)}$ where the kernel is given by
$k_{G}(x,y;d;L_{b},w_{b},\mu_{b})=\frac{1}{2\pi\sigma(d,L_{b},w_{b},\mu_{b})^{2}}\mathrm{exp}\left(-\frac{x^{2}+y^{2}}{2\sigma(d,L_{b},w_{b},\mu_{b})^{2}}\right)$ (6)
$\sigma(d,L_{b},w_{b},\mu_{b})=\frac{1}{2\sqrt{2\log(2)}}\left(\frac{w_{b}}{L_{b}-2/\mu_{b}}\cdot d+w_{b}\right)$ (7)
Convolution with a Gaussian function has computational advantages since it can be decomposed into successive application of two perpendicular 1D kernels via $K_{G}^{(x,y;d;b)}=K_{G}^{(x;d;b)}K_{G}^{(y;d;b)}$; 1D convolution is significantly more computationally efficient than 2D. Unfortunately, the decomposition of a 2D kernel into two 1D kernels is not mathematically possible when the CDR contains significant SPRF and SSRF components. 2D convolution, however, is not the only way to implement Equation 5. Owing to the discrete rotational symmetries and features of the anisotropic PSFs obtained with typical SPECT collimators, we propose the following “1D rotation” model (abbreviated as “1D-R”) for the PSF operator:
$M_{\mathrm{1D-R}}^{(x,y;d;b)}\equiv\left(\sum_{\theta\in\Theta_{t}}\mathcal{R}^{-1}_{\theta}K^{(x;d;b)}_{t}\mathcal{R}_{\theta}+K_{B}^{(x;d;b)}K_{B}^{(y;d;b)}+1\right)K^{(x;d;b)}_{G}K^{(y;d;b)}_{G}$ (8)
where $\mathcal{R}_{\theta}$ is a rotation operator that implements rotation about an axis in the $d$ direction by angle $\theta$, and $\Theta_{t}$ correspond to the angle of the septal penetration tails (equal to $\{0,\pi/3,2\pi/3\}$ for 225Ac). The derived linear operator only makes use of 1D convolutions and rotations, since they often require less computational time on GPU compared to 2D convolutions. The kernels, modeling the components of the CDR (as shown in Figure 1), are selected as follows:
  1. 1.
    Gaussian kernel $K^{(x;d;b)}_{G}$:
    $k_{G}(x;d;b)=A_{G}(d,b)\cdot\mathrm{exp}\left(-x^{2}/2\sigma_{G}(d,b)^{2}\right)$ (9)
    $A_{G}(d,b)=b_{0}e^{-b_{1}d}+b_{2}e^{-b_{3}d}$ (10)
    $\sigma_{G}(d,b)=b_{4}+b_{5}\left(\sqrt{d^{2}+b_{6}^{2}}-|b_{6}|\right)$ (11)
    This kernel is used to build geometric component of the PSF.
  2. 2.
    Tail kernel $K^{(x;d;b)}_{t}$:
    $k_{t}(x;d;b)=A_{t}(d,b)f_{t}(x/\sigma_{t}(d,b))$ (12)
    $A_{t}(d,b)=b_{0}e^{-b_{1}d}+b_{2}e^{-b_{3}d}$ (13)
    $\sigma_{t}(d,b)=1+b_{4}\left(\sqrt{(d-d_{\mathrm{min}})^{2}+b_{5}^{2}}-|b_{5}|\right)$ (14)
    where $d_{\mathrm{min}}$ is the source-detector distance used in the PSF fit. $f_{t}(x)$ is a discrete array of numbers that is linearly interpolated between its fixed points; these points are also considered to be part of the hyperparameters $b$. This kernel is used to build the septal penetration component of the PSF.
  3. 3.
    Isotropic background kernel $K^{(x;d;b)}_{B}$:
    $k_{B}(x;d;b)=A_{B}(d,b)f_{B}(x/\sigma_{B}(d,b))$ (15)
    $A_{B}(d,b)=b_{0}e^{-b_{1}d}+b_{2}e^{-b_{3}d}$ (16)
    $\sigma_{B}(d,b)=1+b_{4}\left(\sqrt{(d-d_{\mathrm{min}})^{2}+b_{5}^{2}}-|b_{5}|\right)$ (17)
    $f_{B}(x)=e^{-|x|}$ (18)
    This kernel is used to build the septal scatter component of the PSF.
Figure 1: 440 keV PSF obtained by simulating a point source at a source-detector distance of 30 cm using SIMIND. Labeled scatter components that refer to the different terms of Equation 8 are shown.
All three model components also encapsulate the small blurring contribution from the IRF. The hyperparameters $b$ are separate for each of the three different components.
In subsequent sections, the proposed 1D-R model is evaluated against a Monte Carlo 2D model where kernel data is obtained using SIMIND [37]; point source projection data is normalized at a large number of source-detector distances. The corresponding PSF operator, abbreviated as “2D”, is defined as
$M_{\mathrm{2D}}^{(x,y;d;b)}\equiv K_{\mathrm{true}}^{x,y;d}$ (19)
where the corresponding kernel $k_{\mathrm{true}}^{x,y;d}$ corresponds to normalized projection data at each source-detector distance $d$.
The publicly SPECTPSFToolbox repository is structured to facilitate the customizability of these models. For the 1D-R model, components are implemented via separate class instances that are subsequently added and multiplied together to obtain the final form of Equation 8. The functional form and hyperparameters of the amplitude/scaling, such as that of Equation 10, are left arbitrary for the user. The 2D model (Equation 19) is implemented via a class instance that receives a stack of PSF kernels with associated source-detector distances: when used in image reconstruction, the model automatically chooses the PSF kernel closest to each source-detector distance in the reconstruction problem.
2.1.2 Image Reconstruction
All acquired image data in subsequent sections are reconstructed using PyTomography with the maximum likelihood expectation maximization (MLEM) algorithm; this was chosen over the ordered subset expectation maximization algorithm (OSEM) because of the low count statistics for each projection. System matrices employed attenuation correction using attenuation maps derived from acquired computed tomography (CT) images. As described by Zekun et al. [38], SPECT with $\alpha$-RPTs often have significant stray radiation-related noise component due to the proportionally low count rate. The imaging system equation is thus given by
$\bar{y}=Hx+\bar{\Psi}+\hat{s}$ (20)
where $\bar{y}$ is the expectation of the acquired count data $y$, $H$ is the system matrix, $x$ is the estimated 3D count distribution, $\bar{\Psi}$ is the expectation of the stray radiation-related noise, and $\hat{s}$ is the scatter estimate. $\bar{\Psi}$ is assumed to be the same for all bins in a given energy window and proportional to the acquisition time, and can be measured by taking a blank scan using the same energy windows as the patient acquisition, averaging the number of counts in each valid bin, and scaling by the acquisition time used in the clinical protocol. Since the dual energy window (DEW) technique is used for scatter estimation, the stray radiation in that window also must be accounted for: the scatter estimation is thus given by
$\hat{s}=\left(\frac{w_{p}}{w_{l}}\right)\left(S_{G}[g_{l}]-\bar{\Psi_{l}}\right)$ (21)
where $w_{p}$ is the width of the primary window, $w_{l}$ is the width of the lower window, $S_{G}$ is a Gaussian smoothing kernel, $g_{l}$ is the acquired count data in the lower energy window, and $\Psi_{l}$ is the expectation of the mean stray-radiation noise in the lower energy window. The smoothing kernel is used here due to the high noise present. One of the dangers of Equation 21 is that it is capable of yielding negative values in Equation 20; in practice this rarely occurred, and all negative values were set to zero.
When used in image reconstruction, the size of the 2D PSF kernels was always set to $N-1\times N-1$ where $N$ is the number of voxels along the largest direction in the reconstructed image. For the 1D-R model, the kernels corresponding to $K_{G}$ and $K_{B}$ were of size $N-1$, while the kernel corresponding to $K_{t}$ was of size $\lceil\sqrt{2}N\rceil+n$ to account for the diagonal, where $n$ is of integer value 0 or 1 to make the kernel size odd. It should be noted that for PSFs much larger than the dimensions of the image, the size of the kernels would need to cover a $12N-1\times 2N-$ 1 area to account for contributions from voxels on one edge of the image to detector elements on the opposite edge; in practice, the PSFs were not that large and the dimensions chosen were sufficient.
2.2 Validation and Timing
A 440 keV point source was simulated using the SIMIND Monte Carlo program [37] at 1100 positions that linearly varied from 0 cm to 58.44 cm. The detector pixel size was $0.24~{}\mathrm{cm}\times 0.24~{}\mathrm{cm}$ with $255\times 255$ pixels. The simulated collimator corresponded to a Siemens high energy configuration with a hexagonal shape, a hole width of 0.40.40.4~{}0.4cm and a hole length of 5.975.975.97~{}5.97cm. The intrinsic resolution of the detector was also included in the simulation and was assumed to be 0.380.380.38~{}0.38cm at 140 keV, representative of a Symbia system with $3/8$” crystal length. The PSFs at all 1100 distances were used to generate $K_{\mathrm{true}}^{x,y;d}$ in Equation 19 during reconstruction by selecting the PSF closest to the actual source-detector distance.
Twelve of these PSFs at source detector distances of 111~{}1cm and every 555~{}5cm from 555~{}5cm to 555555~{}55cm were used as data to fit the 1D-R model. Fitting was performed in three steps: (i) the Gaussian parameters were optimized using the Adaptive Moment Estimation (ADAM) algorithm for $1\cdot 10^{4}$ iterations and a learning rate of $10^{-2}$, (ii) all the parameters were simultaneously optimized using ADAM for $1.5\cdot 10^{4}$ iterations and a learning rate of $10^{-3}$.
The computational time for the 1D-R model (Equation 8) and 2D model (Equation 19) were bench-marked on CPU and GPU for by applying the operators to matrices of four different sizes: $64^{3}$, $128^{3}$, $196^{3}$, and $256^{3}$. Each matrix of size $N^{3}$ was filled with random uniform numbers between 0 and 1. Each experiment used conditions similar to SPECT imaging, where the $d$ axis varied from 000~{}cm to 505050~{}50cm, and the $x$ and $y$ axes varied from $-30.74~{}$cm to 30.7430.7430.74~{}30.74cm. Use of standard convolution and fast fourier transform (FFT) based convolution were compared for each method.
2.3 SPECT Studies
2.3.1 Phantom Study
Application of (i) standard Gaussian (ii) 1D-R (iii) 2D PSF modeling was evaluated for GPU reconstruction of acquired 225Ac data. A cylindrical phantom with spheres of diameter 606060~{}60mm, 282828~{}28mm, and 222222~{}22mm with an initial sphere activity concentration of 1.37 kBq/mL was filled at a 10:1 source to background ratio. 34 SPECT acquisitions of the phantom were taken in sequence on a Symbia T2 SPECT/CT system (Siemens Healthineers, USA) with the following settings: $128\times 128$ pixels at $4.82~{}\mathrm{mm}\times 4.82~{}\mathrm{mm}$ resolution, 96 projection angles, high energy collimators, and 60 s acquisition time per projection. For reconstruction of the 440440440~{}440keV emission from the 213Bi daughter, two energy windows were configured for acquisition: (i) photopeak window centered at 440 keV with a width of 20% and (ii) a lower scatter window used for DEW scatter correction centered at 374374374~{}374keV with a width of 12.5%. A blank scan with identical parameters was acquired to obtain the mean-stray radiation noise in each energy window. Two noise levels of the data were considered:
  1. 1.
    One of the 34 scans was used for data reconstruction. The number of acquired counts in this scenario corresponds to an expected clinical 225Ac scan of a patient undergoing 225Ac-PSMA-617 therapy with injection of 888~{}8MBq [23] and is scanned for 2.5 min per projection at a time point between 0-72 hr post injection.
  2. 2.
    The counts from all 34 scans were summed together and used for image reconstruction. In this case, no scatter blurring was used in Equation 21 since the noise level was low. Because the detectors had an approximately equal radial path around the phantom for each scan, this corresponds to a single scan with a high count rate.
Although 96 projection angles were acquired, only 32 were used since this was similar to the corresponding 225Ac patient acquisition in the next section. All images were reconstructed with MLEM for up to 100 iterations.
2.3.2 Patient Study
A patient receiving 225Ac-PSMA-617 therapy with an injected activity of 8 MBq was imaged 20.5 hours post injection (first cycle). SPECT data were acquired on a Discovery 670 Pro SPECT/CT (GE Healthcare, USA) with a high energy general purpose collimator. 30 projections (15 per head) were acquired for 150 s using an identical photopeak window to the phantom acquisition, and 358.15 keV - 395.85 keV lower window. Since this GE scanner has a different CDR than the Siemens scanner used in prior examples, PSF data were regenerated in SIMIND using a high energy general purpose GE collimator and another 1D-R model was fit as before. In GE scanners, the hexagonal bores in the collimator are rotated by $90^{\circ}$ compared to the Siemens collimator, so the angles in $\Theta_{t}$ were adjusted to compensate for this. The acquired data were reconstructed using (i) standard Gaussian (ii) 2D and (iii) 1D-R PSF modeling using MLEM for up to 100 iterations. Reconstructed images were filtered using a Gaussian filter with a full width half max (FWHM) of 333~{}3cm. Two high uptake bone lesions were segmented on the 100th iteration reconstruction using 3D Slicer [39]. Three lesion ROIs were segmented by a physician on a pre-therapeutical PET image using the PET Edge+ tool of MIM v7.2.1 (MIM Software Inc., USA). A spherical ROI of diameter 101010~{}10cm was placed in a low uptake region in the center of the patient so that the contrast, defined as the mean uptake ratio between the lesion ROIs and background ROI, could be obtained. The mean number of counts and contrast for each lesion ROI were then evaluated for each iteration of MLEM; before the statistics were computed, the 333~{}3cm FWHM Gaussian filter was applied to the image.
3 Results
Figure 2 shows the Monte Carlo simulated PSF data compared to the fit obtained via Equation 8. The curve fit is a reasonable approximation to the PSF at most distances, but is unable to capture features at larger distances, such as the intensity of the tails at 50 cm. The model also consistently under predicts the intensity at the center of the PSF, and this effect is more prevalent at large distances.
Figure 2: Comparison between 225Ac Monte Carlo simulated PSF data and the corresponding fit from the 1D-R model at different source-detector distances. From top to bottom, (i) 2D profiles of Monte Carlo PSF data, (ii) 2D profiles of 1D-R model, (iii) central vertical profile of Monte Carlo and fitted PSFs, (iv) central horizontal profile of Monte Carlo and fitted PSFs; the Monte Carlo data is shown in black scatter points, while the fitted data is shown as a solid red line.
Timing benchmarks for PSF modeling using the 1D-R, 1D-R (FFT), 2D, and 2D (FFT) are shown in Figure 3. CPU implementation yields no benefits with the proposed 1D-R model, and performs fastest with FFT-based 2D convolutions and slowest using regular 2D convolutions. The GPU implementation is faster than the CPU implementation for all methods, and yields computational benefits when the proposed 1D-R model is used. With a matrix size of $128^{3}$, use of the 1D-R method is over three times faster than 2D (FFT) PSF modeling. For small matrix sizes ($64^{3}$) there is no computational speed-up using the 1D-R method, and for large matrix sizes ($256^{3}$) the relative time difference between the 1D-R and 2D (FFT) begins to decrease, approaching only a two times speed advantage.
Figure 3: CPU and GPU required run time of the 1D-R and 2D PSF models for different matrix sizes. Each implementation is evaluated using standard and FFT based convolution. The time shown on top of each bar in units of ms.
PyTomography reconstructed 225Ac phantom images are shown in Figure 4. The time required for 100 iterations of MLEM for the low count data was 110110110~{}110s (Gaussian PSF), 146146146~{}146s (1D-R PSF) and 377377377~{}377s (2D PSF). Reconstructions using the 1D-R and 2D methods are almost indistinguishable qualitatively. All 34 low count acquisitions were then reconstructed and the 1D-R and 2D methods were quantitatively compared; the differences between the 2D and 1D-R were $(1.25\pm 1.25)\%$, $(0.60\pm 0.69)\%$, and $(1.90\pm 0.18)\%$ from smallest to largest sphere respectively. For comparison, the variability in counts between separate noise realizations for the 2D model were $54.4\%$, $33.0\%$, and $9.8\%$ from smallest to largest sphere, respectively.
Figure 4: Axial slices from reconstruction of low-count (top row) and high-count (bottom row) 225Ac phantom data acquired on a Symbia T2 SPECT system. Reconstructed SPECT and CT images are superimposed; the boundaries of each sphere can be seen on the CT. Images were reconstructed with the reference 2D, Gaussian, and proposed 1D-R PSF modeling using the acquired Monte Carlo data. The displayed low count images were post-smoothed using a 3D Gaussian function with a 3 cm FWHM. The plots on the right show the average counts per voxel in the small (sml) medium (med) and large (lrg) spheres as a function of iteration number; the 2D is shown in black, the Gaussian is shown in blue, and the 1D-R is shown in red.
PyTomography reconstructed 225Ac-PSMA-617 patient images are shown in Figure 5. The time required for 100 iterations of MLEM was $38.5$ s (Gaussian PSF), 165.6165.6165.6~{}165.6s (1D-R PSF) and 351.5351.5351.5~{}351.5s (2D PSF). Qualitatively, the Gaussian PSF yields less counts in the uptake ROIs, and also greater background counts, for example, in the vertebrae. The count differences between the 1D-R and 2D methods after 100 iterations were 1.271.271.27~{}1.27(lesion 1), 0.480.480.48~{}0.48(lesion 2) and $2.03~{}\%$ (lesion 3). These differences are comparable to those observed in the phantom study, and were significantly less than the differences between the Gaussian and 2D methods of $-42.3~{}\%$ (lesion 1), $-29.1~{}\%$ (lesion 2) and $-44.3~{}\%$ (lesion 3). The differences in lesion-background contrast between the 1D-R and 2D methods after 100 iterations were $-0.18~{}\%$ (lesion 1), $-0.96~{}\%$ (lesion 2) and $0.56~{}\%$ (lesion 3). These were smaller in magnitude than the contrast differences between the Gaussian and 2D methods of $-51.3\%$ (lesion 1) and $-23.3~{}\%$ (lesion 2) and $-56.7~{}\%$ (lesion 3). Based on the line plots in Figure 5, the Gaussian PSF yields convergence with less iterations; this observation is similar to how an absence CDR modeling leads to faster convergence in traditional low and medium energy SPECT. In general, the smaller the spatial extent of the PSF, the faster convergence is attained in MLEM.
Figure 5: Coronal slices of 68Ga-PSMA-617 (top row, left) and 225Ac-PSMA-617 patient data (top row, right), counts in bone lesion ROIs as a function of iteration number (middle row) and signal to background ratios in the bone lesion ROIs (bottom row). The Background ROI was a 10 cm sphere placed in the center of the patient with little activity. The 68Ga-PSMA-617 image was reconstructed using QClear on the vendor’s software. 225Ac images were reconstructed in PyTomography using (i) reference 2D, (ii) Gaussian, and (iii) the proposed 1D-R model with 100 iterations of MLEM. The displayed reconstructed SPECT images were post filtered using a 3D 3 cm FWHM Gaussian. For computation of the data in the line plots, the image at each iteration was post-filtered using a 3 cm FWHM 3D Gaussian filter before the statistic was computed. Arrows pointing to the three bone lesions are shown in red (lesion 2 is directly on top of its annotation). There was no uptake in the right kidney due to organ failure.
4 Discussion
In this work, a 1D-R PSF model for high energy SPECT reconstruction was proposed, and demonstrated to be both efficacious and fast for use in high-energy reconstruction of 225Ac data. The model was created using publicly shared SPECTPSFToolbox component of the PyTomography project; t is hoped that the shared toolbox and corresponding tutorials will permit the nuclear medicine community to develop custom PSF models for new isotope/collimator configurations as research in novel RPTs and imaging techniques continues to grow.
As shown in Figure 3, the proposed 1D-R only yields computational benefits when image reconstruction is performed on GPU. If reconstruction is performed on CPU, then PSF modeling should be implemented with FFT-based 2D convolutions. FFT-based 2D convolutions are still, however, three times faster on GPU than on CPU. Furthermore, if a GPU is used then the 1D-R model can be used for further computational advantages. The fastest implementation of PSF modeling on GPU (proposed 1D-R) is approximately one order of magnitude faster than the fastest implementation of PSF modeling on CPU (FFT-based 2D convolutions).
While the proposed 1D-R model is not able to perfectly capture the features of the PSF at all radial distances in Figure 2, it yields near identical reconstructions in Figures 4 and 5, especially when compared to reconstruction using the Gaussian kernel. In the low count scenario of the phantom study, which approximately represents a clinical count rate, the percent difference in the three spheres between the 1D-R and 2D model were small compared to the variability between separate acquisitions. In the patient example, the variability of counts between the 2D and 1D-R models in the bone lesions was similar to the spheres in the phantom: this suggests that any differences in predicted counts between the 2D and 1D-R models in the patient example is also negligible compared to differences that would be observed between separate acquisitions. While the 1D-R model and 2D model yielded similar contrast, the Gaussian model had significantly reduced contrast with differences of $-51.3~{}\%$, $-23.3~{}\%$, and $-56.7~{}\%$ compared to the 2D model in the bone lesions. While scaling the magnitude of the image reconstructed using the Gaussian PSF could artificially increase the number of counts in each bone ROI, it would have no impact on these substantial differences in contrast.
In the phantom example, the radial path of the detector approached 7.27.27.2~{}7.2cm at its minimum and 27.227.227.2~{}27.2cm at its maximum. Since the radius of the cylinder was 101010~{}10cm, this suggests that the PSF model was able to reasonably function between $0-37~{}$cm. Future work may seek to test the efficacy of the proposed 1D-R model for larger radial distances, as this may be required for larger patients.
While the functional form of Equations 9, 12, 15 were selected because they produced a reasonable approximation of the PSF, it may be the case that they can be improved upon. Use of rotation operations with the component $K_{B}^{(x;d;b)}K_{B}^{(y;d;b)}$ may permit more radial symmetry in the SSRF. Substitution of $K_{G}^{(x;d;b)}K_{G}^{(y;d;b)}$ with a small 2D kernel that represents the true aperture function of the GRF may remedy the under prediction in the center of the PSFs at large distances. Close inspection of the SIMIND data at large distances in Figure 2 reveals additional dim tails; these could be modeled by including another independent tail component in the model. A trade-off for adding additional features to the model, however, is an increase in computation time. The SPECTPSFToolbox Python library released along with this paper contains tutorials demonstrating how to create custom and fit parameterized PSF models. Users can obtain point source data using Monte Carlo programs (such as GATE [40] and SIMIND) or by using real scanners, and fit corresponding models to the acquired data. The models can then be imported to PyTomography for customized and scanner specific SPECT reconstruction. Users can independently evaluate the trade-off between adding model features, impact on reconstructed images, and increase in computation time.
As discussed in Section 2.1.2, the kernel sizes were fixed to $N-1\times N-1$ for all $d$. For small $d$, where the PSF is also smaller, a kernel of this size might be excessive and could unnecessarily increase the computational time required. Future research exploring the reduction of computational time might experiment with using a source-detector distance dependent kernel size $N(d)$ for that matches the size of the PSF; this would further reduce the computational time required.
Since reconstruction using the proposed 1D-R model more than halves reconstruction times compared to the 2D model, it may be preferable for reconstruction of 225Ac data. Both the 1D-R and 2D models, however, are only applicable when Equation 5 is used for SPECT system matrix modeling; this equation relies on the assumption that the attenuation probabilities vary little across the PSF. Due to the large spatial extent of the PSFs shown here, this assumption may be invalid, and Equation 4 may instead be required for accurate reconstruction of 225Ac data. This assumption largely depends on the density of the object being scanned. If the object had a region of high density which only rays on the outer edges of the PSF passed through, then the equal attenuation path assumption would be invalid. For patients and phantoms, where the density is roughly constant throughout the field of view, application of Equation 5 may be permitted. Meanwhile, the study of different density configurations and the applicability of Equation 5 remains to be investigated in future work.
In conclusion, a fast and accurate implementation of high energy CDR modeling in SPECT imaging that uses 1D convolutions and rotations was developed, and tested on 225Ac reconstructions. The technique was implemented using the open source reconstruction library PyTomography, and was shown to speed up reconstruction times by more than a factor of two compared to conventional, 2D convolution based methods. Furthermore, the 1D-R method yielded near identical results to the 2D method, and the small differences were insignificant compared to the differences between reconstructions of separate noise realizations. The SPECTPSFToolbox python library was developed and publicly shared to permit others in the nuclear medicine community to develop custom isotope/collimator PSF models for use in the open-source reconstruction library PyTomography.
5 Acknowledgements
This work was supported by the Natural Sciences and Engineering Research Council of Canada (NSERC) CGS D Award 569711 and Discovery Grant RGPIN-2019-06467, as well as computational resources and services provided by Microsoft AI for Health.

### v2

@@
-Fast and Accurate Collimator-Detector Response Compensation in High-Energy SPECT Imaging with 1D Convolutions and Rotations
+Computationally Efficient Collimator-Detector Response Compensation in High Energy SPECT using 1D Convolutions and Rotations
 Abstract
-Objective: Modeling of the collimator-detector response (CDR) in SPECT reconstruction enables improved resolution and more accurate quantitation, especially for higher energy imaging (e.g.
-177Lu and 225Ac).
-Such modeling, however, can pose a significant computational bottleneck when there are substantial components of septal penetration and scatter in the acquired data, since a direct convolution-based approach requires large 2D kernels.
-The present work presents an alternative method for fast and accurate CDR compensation using a linear operator built from 1D convolutions and rotations (1D-R).
+Objective: Modeling of the collimator-detector response (CDR) in SPECT reconstruction enables improved resolution and accuracy, and is thus important for quantitative imaging applications such as dosimetry.
+The implementation of CDR modeling, however, can become a computational bottleneck when there are substantial components of septal penetration and scatter in the acquired data, since a direct convolution-based approach requires large 2D kernels.
+This work proposes a 1D convolution and rotation-based CDR model that reduces reconstruction times but maintains consistency with models that employ 2D convolutions.
 To enable open-source development and use of these models in image reconstruction, we release a SPECTPSFToolbox repository for the PyTomography project on GitHub.
-Approach: A 1D-R CDR model was formulated, and subsequently fit to Monte Carlo 440 keV point source data representative of emissions in 225Ac imaging.
-Computation times of (i) the proposed 1D-R model and (ii) a traditional model that uses 2 dimensional convolutions (2D) were compared for typical SPECT matrix sizes.
-Both CDR modeling techniques were then used to reconstruct 225Ac phantom and patient data, and were compared by quantifying total counts in hot regions of interest (ROIs) and activity contrast between hot ROIs and background regions.
-Results: The 1D-R and 2D CDR models were created using the SPECTPSFToolbox.
-For typical matrix sizes in SPECT reconstruction, application of the 1D-R model provides a two-fold computational speed-up over the 2D model.
-Only small differences between the 1D-R and 2D models (order of $1\%$ ) were obtained for count and contrast quantification in select ROIs.
+Approach: A 1D/rotation-based CDR model was formulated and subsequently fit to Monte Carlo point source data representative of 177Lu, 131I, and 225Ac imaging.
+Computation times of (i) the proposed 1D/rotation-based model and (ii) a traditional model that uses 2D convolutions were compared for typical SPECT matrix sizes.
+Both CDR models were then used in the reconstruction of Monte Carlo, physical phantom, and patient data; the models were compared by quantifying total counts in hot regions of interest (ROIs) and activity contrast between hot ROIs and background regions.
+Results: For typical matrix sizes in SPECT reconstruction, application of the 1D/rotation-based model provides a two-fold computational speed-up over the 2D model when running on GPU.
+Only small differences between the 1D/rotation-based and 2D models (order of $1\%$) were obtained for count and contrast quantification in select ROIs.
 Significance: A technique for CDR modeling in SPECT was proposed that (i) significantly speeds up reconstruction times, and (ii) yields nearly identical reconstructions to traditional 2D convolution based CDR techniques.
@@
 1 Introduction
-Single photon emission computed tomography (SPECT) imaging is an in vivo modality of significant value in different clinical applications [1, 2, 3].
-In particular, quantitative SPECT imaging holds great potential and value in the field of theranostics, wherein image quantification and dosimetry can be performed to improve radiopharmaceutical therapies (RPTs).
+Single photon emission computed tomography (SPECT) is an in vivo imaging modality, possessing value for various clinical applications [1, 2, 3].
+In particular, quantitative SPECT imaging has applications in the field of theranostics, wherein image quantification and dosimetry can be performed to improve radiopharmaceutical therapies (RPTs).
 Routine SPECT-based dosimetry could unveil the relationships between tumor-response and healthy organ complications with absorbed dose, therefore enabling personalized treatments that maximize tumor dose while minimizing toxicity to organs at risk [4, 5].
@@
 The system model is a linear operator that predicts the expectation of the acquired data given a 3D isotope distribution, accounting for phenomena such as photon attenuation in the patient.
-An important aspect of the model is collimator detector response (CDR) modeling, which estimates image blurring caused by the collimator and the scintillation crystals [10, 11, 12].
+An important aspect of the model is collimator detector response (CDR) modeling, which estimates image blurring caused by the collimator and the detector [10, 11, 12].
 Computation of the CDR is typically the most computationally expensive operation in image reconstruction.
-The CDR can be characterized by the detector response from a point source of activity; this is denoted the point spread function (PSF).
+The CDR can be characterized by the system’s response to a point source of activity; this is denoted the point spread function (PSF).
 The PSF can be decomposed into multiple components.
-The intrinsic response function (IRF) characterizes the inability of the scintillation crystals to precisely localize the point of interaction, and is sufficiently modeled using a Gaussian function.
-The collimator response for parallel hole collimators results from the inability of the collimator to accept only photons travelling perpendicular to the detector; it consists of three components: (i) the geometric response function (GRF) [11, 13, 14, 15], which describes photons that travel through the collimator holes without penetrating or interacting with the septa, (ii) the septal penetration response function (SPRF), which describes the contribution from photons that travel through the collimator without being attenuated, and (iii) the septal scatter response function (SSRF), which consists of photons that interacted and scattered within the collimator and were subsequently detected in the scintillator.
-Selection of collimator parameters is an important aspect of SPECT imaging [16].
-As the collimator becomes thicker and the diameter of the collimator bores becomes narrower, the relative contribution from the SPRF and SSRF decreases due to the increased attenuation probability for photons not travelling perpendicular to the detector.
-In this case, (i) the detector resolution improves and (ii) the net point spread function (PSF) is dominated by the IRF + GRF and can be reasonably approximated using a 2D Gaussian function.
-A trade-off of having a collimator with high septal thickness and small hole diameters, however, is a decrease in detector sensitivity; the corresponding implications for quantitative imaging is a decrease in precision or longer patient scan times.
+The intrinsic response function (IRF) characterizes the uncertainty on the point of interaction within the detecting material (e.g.
+an NaI scintillator), and is sufficiently modeled using a Gaussian function.
+The collimator response for parallel hole collimators results from the inability of the collimator to accept only photons travelling perpendicular to the detector; it consists of three components: (i) the geometric response function (GRF) [11, 13, 14, 15], which describes photons that travel through the collimator holes without penetrating or interacting with the septa, (ii) the septal penetration response function (SPRF), which describes the contribution from photons that travel through the collimator without being attenuated, and (iii) the septal scatter response function (SSRF), which consists of photons that interacted and scattered within the collimator and were subsequently detected.
+The selection of collimator parameters is an important aspect of SPECT imaging [16].
+A collimator with longer, narrower bores, and thicker septal thickness reduces the contribution from the SPRF and SSRF by increasing the probability of attenuation for photons not traveling perpendicular to the detector.
+If the collimator geometry is optimized as such, the net point spread function (PSF) is dominated by the IRF + GRF and can be reasonably approximated using a 2D Gaussian function.
+A trade-off of having a thicker collimator with smaller hole diameters, however, is a decrease in detector sensitivity; the corresponding implications for quantitative imaging is a decrease in precision or longer patient scan times.
 This trade-off must be independently considered for each isotope.
-For 177Lu labeled radiopharmaceuticals, where the imaged photons are 208208208~{}208keV, a commercially labeled “medium energy” collimator configuration (i) yields a reasonable count rate and (ii) adequately minimizes the SPRF and SSRF components.
+For 177Lu labeled radiopharmaceuticals, where the 208208208~{}208keV photopeak is typically used for imaging, a commercially labeled “medium energy” collimator configuration (i) yields a reasonable count rate and (ii) adequately minimizes the SPRF and SSRF components.
 In this situation, a 2D Gaussian PSF model is sufficient for image reconstruction [17, 18].
-Many commercially available reconstruction software only offer Gaussian PSF modeling, and thus implicitly assume the SPRF and SSRF are negligible.
+To the knowledge of the authors, all presently available commercial reconstruction software offer Gaussian PSF modeling, and thus implicitly assume the SPRF and SSRF are negligible.
 There are multiple advantages to this.
@@
 This is typically an issue with radioisotopes that emit high photon energies, such as $\alpha$-emitters like 225Ac.
-Use of $\alpha$-emitters in radiopharmaceutical therapies presents a major and exciting frontier, due to the high linear energy transfer (LET) associated with $\alpha$ emissions [19, 20, 21, 22, 23, 24].
+Use of $\alpha$-emitters in radiopharmaceutical therapies presents an exciting frontier due to the high linear energy transfer (LET) associated with $\alpha$ particles, which may yield benefits for tumor eradication [19, 20, 21, 22, 23, 24].
 In preclinical studies, simultaneous treatment with 177Lu-PSMA-617 and 225Ac-PSMA-617 compared to 177Lu-PSMA-617 alone resulted in significantly reduced tumor growth [25].
@@
 [35] recently showed that the limited-memory Broyden-Feltcher-Goldfarb-Shannon algorithm with box constraints and a diagonal preconditioner (L-BFGS-B-PC) was able to converge several times faster than the one step late expectation maximum (OSL-EM) algorithm in Positron Emission Tomography (PET) reconstruction because less projection operations were required.
-As another example, in Chun et al.
-[34], long reconstruction times for PSF modeling of 131I were partially remedied by using fast fourier transform (FFT) based 2D convolutions.
+As another example, Chun et al.
+[34] reduced reconstruction times of 131I by using fast fourier transform (FFT) based 2D convolutions for PSF modeling.
 By contrast, in the present work, we seek to reduce reconstruction times by using a PSF model that incorporates (i) 1D convolutions and (ii) rotations.
-It will be shown that this model significantly reduces computational time compared to standard and FFT based 2D convolutions, and consequently significantly reduces the time required for SPECT image reconstruction of 225Ac.
-The method is evaluated on phantom data and patient data, and we show that the proposed model yields identical results to Monte Carlo based 2D convolution techniques.
-All models considered in this paper are implemented using the GPU-accelerated functionality of PyTorch.
+It will be shown that this model runs faster than standard and FFT based 2D convolutions when implemented on a GPU, and can thus be used to speed up image reconstruction times.
+The method is evaluated on Monte Carlo phantom data of 177Lu, 131I, and 225Ac, as well as physical phantom and patient data of 225Ac.
+In each case, the method demonstrates near-identical recovery coefficients and contrast in regions of interest compared to MC-based 2D convolution techniques.
 Alongside this paper, we release the SPECTPSFToolbox: an open-source GitHub repository which forms a new component of the PyTomography [36] project.
@@
 2 Materials and Methods
-In Section 2.1, the mathematical formalism of the PSF model is outlined, and the reconstruction protocol for all acquired data in subsequent sections is established.
-Section 2.2 describes how the model is fit to SIMIND [37] Monte Carlo 225Ac point source data acquisitions simulated at various source-detector distances.
+Section 2.1 outlines the mathematical formalism of the PSF model and establishes the reconstruction protocol for all acquired data presented in subsequent sections.
+Section 2.2 describes how the model is fit to Monte Carlo (MC) 177Lu, 131I, and 225Ac point source data acquisitions simulated at various source-detector distances.
 The computational time of the model is benchmarked and compared to 2D PSF modeling and Gaussian PSF modeling.
-In Section 2.3, the model is then used for reconstruction of (i) an 225Ac phantom consisting of spherical inserts and (ii) a patient receiving 225Ac-PSMA-617 treatment for metastatic prostate cancer.
+In Section 2.3, the model is then used for reconstruction of (i) MC simulated 177Lu, 131I, and 225Ac phantoms, (ii) a physical 225Ac phantom, and (iii) a patient receiving 225Ac-PSMA-617 treatment for metastatic prostate cancer.
 All computation was performed using a Microsoft Azure virtual machine (Standard NC6s v3) with a 6 CPUs (Intel(R) Xeon(R) CPU E5-2690 v4 @ 2.60GHz), 112 GB of RAM, and a Tesla V100 GPU.
+All MC data was generated using SIMIND [37].
+In order to capture the effects of septal penetration and scatter, SIMIND has an optional routine based on the delta-scattering technique [38], which has been validated for high-energy isotopes such as 131I [39].
+This SIMIND routine has been used in previous studies to generate stacks of PSFs at various source-detector distances for MC-based 225Ac PSF modeling in custom reconstruction [23], since the reconstruction software of the scanner manufacturers does not provide comprehensive PSF modeling of high energy isotopes.
 2.1 Theory
@@
 $K^{(x;d;b)}f\equiv\sum_{x^{\prime}}f(x^{\prime},y;d)k(x-x^{\prime};d;b)$ (1)
-$K^{(y;d;b)}f\equiv\sum_{y^{\prime}}f(x,y^{\prime};d)k(y-y^{\prime};d)$ (2)
+$K^{(y;d;b)}f\equiv\sum_{y^{\prime}}f(x,y^{\prime};d)k(y-y^{\prime};d;b)$ (2)
 $K^{(x,y;d;b)}f\equiv\sum_{x^{\prime},y^{\prime}}f(x^{\prime},y^{\prime};d;b)k(x-x^{\prime},y-y^{\prime};d;b)$ (3)
@@
 where $(x^{\prime},y^{\prime})$ is the source position on a plane parallel to the detector in 3D space, $\mathrm{psf}(...,d)$ is a kernel that yields the point spread function (PSF) at a distance $d$ from the detector, and $p_{\mathrm{att}}(x,y,x^{\prime},y^{\prime},d)$ is the probability that photons traveling from $(x,y,d)$ to detector coordinate $(x^{\prime},y^{\prime})$ are not attenuated.
-Under the assumption that the attenuation probabilities vary little across the PSF, it follows that $p_{\mathrm{att}}(x,y,x^{\prime},y^{\prime},d)\approx p_{\mathrm{att}}(x,y,x,y,d)$ and it can simply be expressed as $p_{\mathrm{att}}(x,y,d)$.
-Defining $f^{\prime}\equiv p_{\mathrm{att}}f$ as the “attenuation-adjusted” image, Equation 4 can then be rewritten in operator form (Equation 3) as
+If the attenuation probabilities are approximately constant for different photon impingement angles permitted on the detector (e.g.
+valid when the PSF is small), it follows that $p_{\mathrm{att}}(x,y,x^{\prime},y^{\prime},d)$ can be replaced with $p_{\mathrm{att}}(x,y,x,y,d)$, which can simply be written as $p_{\mathrm{att}}(x,y,d)$.
+Under this assumption, $f^{\prime}\equiv p_{\mathrm{att}}f$ can be defined as the “attenuation-adjusted” image, and Equation 4 can then be rewritten in operator form (Equation 3) as
 $g_{\phi}(x,y)=\sum_{d}K_{\mathrm{PSF}}^{(x,y;d;b)}f^{\prime}$ (5)
-Since this convolution operation is typically the bottleneck of SPECT system matrix modeling and image reconstruction, it is of interest to look for techniques to reduce the computation time.
-Under conditions of no septal penetration and scatter, the CDR is dominated by the GRF: $K_{\mathrm{PSF}}^{(x,y;d;b)}$ can then be sufficiently approximated using a 2D Gaussian convolution $K_{G}^{(x,y;d;b)}$ where the kernel is given by
+Since this convolution operation often forms the bottleneck of SPECT system matrix modeling and image reconstruction, it is of interest to look for techniques to reduce the computation time.
+Under conditions of no septal penetration and scatter, the CDR is dominated by the GRF: $K_{\mathrm{PSF}}^{(x,y;d;b)}$ can then be sufficiently approximated using a 2D Gaussian convolution $K_{\mathrm{2DG}}^{(x,y;d;b)}$ where the kernel is given by
 $k_{G}(x,y;d;L_{b},w_{b},\mu_{b})=\frac{1}{2\pi\sigma(d,L_{b},w_{b},\mu_{b})^{2}}\mathrm{exp}\left(-\frac{x^{2}+y^{2}}{2\sigma(d,L_{b},w_{b},\mu_{b})^{2}}\right)$ (6)
 $\sigma(d,L_{b},w_{b},\mu_{b})=\frac{1}{2\sqrt{2\log(2)}}\left(\frac{w_{b}}{L_{b}-2/\mu_{b}}\cdot d+w_{b}\right)$ (7)
-Convolution with a Gaussian function has computational advantages since it can be decomposed into successive application of two perpendicular 1D kernels via $K_{G}^{(x,y;d;b)}=K_{G}^{(x;d;b)}K_{G}^{(y;d;b)}$; 1D convolution is significantly more computationally efficient than 2D.
+Convolution with a Gaussian function has computational advantages since it can be decomposed into successive application of two perpendicular 1D kernels via $K_{\mathrm{2DG}}^{(x,y;d;b)}=K_{\mathrm{1DG}}^{(x;d;b)}K_{\mathrm{1DG}}^{(y;d;b)}$; 1D convolution is significantly more computationally efficient than 2D.
 Unfortunately, the decomposition of a 2D kernel into two 1D kernels is not mathematically possible when the CDR contains significant SPRF and SSRF components.
 2D convolution, however, is not the only way to implement Equation 5.
-Owing to the discrete rotational symmetries and features of the anisotropic PSFs obtained with typical SPECT collimators, we propose the following “1D rotation” model (abbreviated as “1D-R”) for the PSF operator:
-$M_{\mathrm{1D-R}}^{(x,y;d;b)}\equiv\left(\sum_{\theta\in\Theta_{t}}\mathcal{R}^{-1}_{\theta}K^{(x;d;b)}_{t}\mathcal{R}_{\theta}+K_{B}^{(x;d;b)}K_{B}^{(y;d;b)}+1\right)K^{(x;d;b)}_{G}K^{(y;d;b)}_{G}$ (8)
-where $\mathcal{R}_{\theta}$ is a rotation operator that implements rotation about an axis in the $d$ direction by angle $\theta$, and $\Theta_{t}$ correspond to the angle of the septal penetration tails (equal to $\{0,\pi/3,2\pi/3\}$ for 225Ac).
-The derived linear operator only makes use of 1D convolutions and rotations, since they often require less computational time on GPU compared to 2D convolutions.
+Owing to the discrete rotational symmetries and features of the anisotropic PSFs obtained with parallel hole SPECT collimators, we propose the following “1D rotation” model (abbreviated as “1DR”) for the PSF operator:
+$M_{\mathrm{1DR}}^{(x,y;d;b)}\equiv\left(\sum_{\theta\in\Theta_{t}}\mathcal{R}^{-1}_{\theta}K^{(x;d;b)}_{t}\mathcal{R}_{\theta}+K_{B}^{(x;d;b)}K_{B}^{(y;d;b)}+1\right)K^{(x;d;b)}_{G}K^{(y;d;b)}_{G}$ (8)
+where $\mathcal{R}_{\theta}$ is a rotation operator that implements rotation about an axis in the $d$ direction by angle $\theta$, and $\Theta_{t}$ correspond to the angle of the septal penetration tails (equal to $\{0,\pi/3,2\pi/3\}$ for hexagonal collimators).
+This derived linear operator only makes use of 1D convolutions and rotations, since they require less computational time on GPU compared to 2D convolutions.
 The kernels, modeling the components of the CDR (as shown in Figure 1), are selected as follows:
@@
     This kernel is used to build geometric component of the PSF.
+The terms $A_{G}(d,b)$ and $\sigma_{G}(d,b)$ are able to capture the decaying amplitude and spreading out of the PSF at larger distances $d$.
   2. 2.
@@
     $k_{t}(x;d;b)=A_{t}(d,b)f_{t}(x/\sigma_{t}(d,b))$ (12)
-    $A_{t}(d,b)=b_{0}e^{-b_{1}d}+b_{2}e^{-b_{3}d}$ (13)
-    $\sigma_{t}(d,b)=1+b_{4}\left(\sqrt{(d-d_{\mathrm{min}})^{2}+b_{5}^{2}}-|b_{5}|\right)$ (14)
+    $A_{t}(d,b)=b_{7}e^{-b_{8}d}+b_{9}e^{-b_{10}d}$ (13)
+    $\sigma_{t}(d,b)=1+b_{11}\left(\sqrt{(d-d_{\mathrm{min}})^{2}+b_{12}^{2}}-|b_{12}|\right)$ (14)
     where $d_{\mathrm{min}}$ is the source-detector distance used in the PSF fit.
-$f_{t}(x)$ is a discrete array of numbers that is linearly interpolated between its fixed points; these points are also considered to be part of the hyperparameters $b$.
-This kernel is used to build the septal penetration component of the PSF.
+$f_{t}(x)$ is a discrete array of numbers that is linearly interpolated between its fixed points.
+This kernel models the tails of the septal penetration PSF component.
+Similar to the geometric component, the terms $A_{t}(d,b)$ and $\sigma_{t}(d,b)$ are such that the amplitude decays and the component spreads out at larger distances $d$.
+The function $f_{t}(x/\sigma_{t}(d,b))$ is initialized as a mono-exponential function to match the decay of the tails observed in Figure 2.
   3. 3.
@@
     $k_{B}(x;d;b)=A_{B}(d,b)f_{B}(x/\sigma_{B}(d,b))$ (15)
-    $A_{B}(d,b)=b_{0}e^{-b_{1}d}+b_{2}e^{-b_{3}d}$ (16)
-    $\sigma_{B}(d,b)=1+b_{4}\left(\sqrt{(d-d_{\mathrm{min}})^{2}+b_{5}^{2}}-|b_{5}|\right)$ (17)
+    $A_{B}(d,b)=b_{13}e^{-b_{14}d}+b_{15}e^{-b_{16}d}$ (16)
+    $\sigma_{B}(d,b)=1+b_{17}\left(\sqrt{(d-d_{\mathrm{min}})^{2}+b_{18}^{2}}-|b_{18}|\right)$ (17)
     $f_{B}(x)=e^{-|x|}$ (18)
     This kernel is used to build the septal scatter component of the PSF.
-Figure 1: 440 keV PSF obtained by simulating a point source at a source-detector distance of 30 cm using SIMIND. Labeled scatter components that refer to the different terms of Equation 8 are shown.
+Similar to the previous two components, the terms $A_{B}(d,b)$ and $\sigma_{B}(d,b)$ capture the amplitude decay and spreading out of the component at larger distances $d$.
+The function form of $f_{B}(x)$ was selected based on the approximate decay of septal scatter seen in Figure 2.
+Figure 1: (a) Top view of collimator, where $w_{b}$ is the collimator hole width.
+(b) Side view of collimator, where $L_{b}$ is the length of the collimator.
+Three different photon detection paths are shown as black arrows; from left to right they correspond to septal scatter, septal penetration, and geometric component respectively.
+(c) 440 keV PSF obtained by simulating a point source at a source-detector distance of 30 cm using SIMIND.
+Labeled scatter components that refer to the different terms of Equation 8 are shown.
 All three model components also encapsulate the small blurring contribution from the IRF. The hyperparameters $b$ are separate for each of the three different components.
-In subsequent sections, the proposed 1D-R model is evaluated against a Monte Carlo 2D model where kernel data is obtained using SIMIND [37]; point source projection data is normalized at a large number of source-detector distances.
-The corresponding PSF operator, abbreviated as “2D”, is defined as
-$M_{\mathrm{2D}}^{(x,y;d;b)}\equiv K_{\mathrm{true}}^{x,y;d}$ (19)
-where the corresponding kernel $k_{\mathrm{true}}^{x,y;d}$ corresponds to normalized projection data at each source-detector distance $d$.
-The publicly SPECTPSFToolbox repository is structured to facilitate the customizability of these models.
-For the 1D-R model, components are implemented via separate class instances that are subsequently added and multiplied together to obtain the final form of Equation 8.
-The functional form and hyperparameters of the amplitude/scaling, such as that of Equation 10, are left arbitrary for the user.
-The 2D model (Equation 19) is implemented via a class instance that receives a stack of PSF kernels with associated source-detector distances: when used in image reconstruction, the model automatically chooses the PSF kernel closest to each source-detector distance in the reconstruction problem.
+PSF modeling can also be accomplished using a generic 2D kernel; the model can be expressed as
+$M_{\mathrm{2D}}^{(x,y;d;b)}\equiv K_{\mathrm{2D}}^{(x,y;d)}$ (19)
+where the corresponding kernel $k_{\mathrm{2D}}^{(x,y;d)}$ might be obtained from real data measurements, MC measurements, or an analytical model based on measurements (such as in [35]).
+In subsequent sections, the following three methods are compared:
+  1. 1.
+    The 2D Gaussian model of Equation 6, with parameters analytically obtained from collimator dimensions. Denoted “2DG”, this is the technique used by available commercial reconstruction software.
+  2. 2.
+    A generic 2D Monte Carlo model (Equation 19) obtained directly from a stack of normalized point source MC data at various source-detector distances.
+Denoted “2D-MC”, this is the same technique used by Delker et.
+al.
+[23] for reconstruction of 225Ac patient data; it will be used as a standard against which method (iii) is compared.
+  3. 3.
+    The proposed 1DR model (Equation 8) with parameters obtained via fitting to MC point source data, and denoted “1DR-MC”.
+Since it is fit to the data used for method (ii), an ideal model would yield identical reconstructions to those obtained with method (ii).
+It should be emphasized that method (i) does not require any empirical data, while methods (ii) and (iii) use MC point source data (hence the inclusion of “-MC” in the acronyms).
+Methods (ii) and (iii), however, can also be implemented using real point source measurements if that data is available.
+The publicly available SPECTPSFToolbox repository is structured to facilitate the customizability of these models.
+For the 1DR-MC model, components are implemented via separate class instances that are subsequently added and multiplied together to obtain the final form of Equation 8.
+The functional form and hyperparameters of the amplitude/scaling, (e.g.
+Equation 10) are left arbitrary for the user; this permits full customization of PSF models.
+The 2D-MC model is implemented via a class instance that receives a stack of PSF kernels with associated source-detector distances: when used in image reconstruction, the model automatically chooses the PSF kernel closest to each source-detector distance in the reconstruction problem.
 2.1.2 Image Reconstruction
-All acquired image data in subsequent sections are reconstructed using PyTomography with the maximum likelihood expectation maximization (MLEM) algorithm; this was chosen over the ordered subset expectation maximization algorithm (OSEM) because of the low count statistics for each projection.
-System matrices employed attenuation correction using attenuation maps derived from acquired computed tomography (CT) images.
-As described by Zekun et al.
-[38], SPECT with $\alpha$-RPTs often have significant stray radiation-related noise component due to the proportionally low count rate.
-The imaging system equation is thus given by
-$\bar{y}=Hx+\bar{\Psi}+\hat{s}$ (20)
-where $\bar{y}$ is the expectation of the acquired count data $y$, $H$ is the system matrix, $x$ is the estimated 3D count distribution, $\bar{\Psi}$ is the expectation of the stray radiation-related noise, and $\hat{s}$ is the scatter estimate.
-$\bar{\Psi}$ is assumed to be the same for all bins in a given energy window and proportional to the acquisition time, and can be measured by taking a blank scan using the same energy windows as the patient acquisition, averaging the number of counts in each valid bin, and scaling by the acquisition time used in the clinical protocol.
-Since the dual energy window (DEW) technique is used for scatter estimation, the stray radiation in that window also must be accounted for: the scatter estimation is thus given by
-$\hat{s}=\left(\frac{w_{p}}{w_{l}}\right)\left(S_{G}[g_{l}]-\bar{\Psi_{l}}\right)$ (21)
-where $w_{p}$ is the width of the primary window, $w_{l}$ is the width of the lower window, $S_{G}$ is a Gaussian smoothing kernel, $g_{l}$ is the acquired count data in the lower energy window, and $\Psi_{l}$ is the expectation of the mean stray-radiation noise in the lower energy window.
-The smoothing kernel is used here due to the high noise present.
-One of the dangers of Equation 21 is that it is capable of yielding negative values in Equation 20; in practice this rarely occurred, and all negative values were set to zero.
-When used in image reconstruction, the size of the 2D PSF kernels was always set to $N-1\times N-1$ where $N$ is the number of voxels along the largest direction in the reconstructed image.
-For the 1D-R model, the kernels corresponding to $K_{G}$ and $K_{B}$ were of size $N-1$, while the kernel corresponding to $K_{t}$ was of size $\lceil\sqrt{2}N\rceil+n$ to account for the diagonal, where $n$ is of integer value 0 or 1 to make the kernel size odd.
-It should be noted that for PSFs much larger than the dimensions of the image, the size of the kernels would need to cover a $12N-1\times 2N-$ 1 area to account for contributions from voxels on one edge of the image to detector elements on the opposite edge; in practice, the PSFs were not that large and the dimensions chosen were sufficient.
+All acquired image data in subsequent sections are reconstructed using PyTomography with the ordered subset expectation maximization algorithm (OSEM); when one subset is used, the algorithm is denoted as maximum likelihood expectation maximization (MLEM).
+System matrices employed attenuation correction using attenuation maps derived from either (i) ground truth attenuation maps from MC simulations or (ii) acquired computed tomography (CT) images in physical studies.
+The imaging system equation is given by
+$\bar{y}=Hx+\hat{s}$ (20)
+where $\bar{y}$ is the expectation of the acquired count data $y$, $H$ is the system matrix, and $x$ is the estimated 3D count distribution.
+$\hat{s}$ is an additive term that accounts for phantom or patient scatter, events that back-scattered in the detector from higher energy emission peaks, and stray background radiation; the presence of stray background radiation for low count SPECT has been described by Zekun et al.
+[40].
+For all cases in this paper, the triple energy window (TEW) technique is used to estimate $\hat{s}$.
+When used in image reconstruction, the size of the 2D PSF kernels in the 2D-MC method was always set to $N-1\times N-1$ where $N$ is the number of voxels along the largest direction in the reconstructed image.
+For the 1DR-MC method, the kernels corresponding to $K_{G}$ and $K_{B}$ were of size $N-1$, while the kernel corresponding to $K_{t}$ was of size $\lceil\sqrt{2}N\rceil+n$ to account for the diagonal, where $n$ is of integer value 0 or 1 to make the kernel size odd.
+It should be noted that for PSFs much larger than the dimensions of the image, the size of the kernels would need to cover a $12N-1\times 2N-$ 1 area to account for contributions from voxels on one corner of the image to detector elements on the opposite corner; in practice, the PSFs were not that large and the dimensions chosen were sufficient.
+2.1.3 Image Quality Metrics
+To compare images reconstructed using the three different PSF models, two image metrics are used. The first metric is recovery coefficients (RCs), defined as
+$\mathrm{RC}\equiv\frac{A_{\mathrm{pred}}}{A_{\mathrm{true}}}$ (21)
+where $A_{\mathrm{pred}}$ corresponds to the predicted activity concentration in a region of interest (ROI), and $A_{\mathrm{true}}$ corresponds to the true activity in that region.
+In MC simulations consisting of spheres with infinite resolution, the ROIs on the reconstructed voxel grid are segmented such that voxels inside ROI are assigned 1 while voxels outside are assigned 0.
+The calibration factors necessary to convert counts to activity units are obtained via point source simulations with the same detector parameters.
+For the physical data, where calibration factors are not available, average counts are instead reported.
+The second metric used is contrast, defined as
+$\mathrm{Contrast}\equiv\frac{C_{\mathrm{hot}}}{C_{\mathrm{bkg}}}$ (22)
+where $C_{\mathrm{hot}}$ is the number of counts in a hot ROI and $C_{\mathrm{bkg}}$ is the number of counts in the background.
 2.2 Validation and Timing
-A 440 keV point source was simulated using the SIMIND Monte Carlo program [37] at 1100 positions that linearly varied from 0 cm to 58.44 cm.
+Point sources with energies of 208 keV, 364 keV, and 440 keV (representative of the typical primary energies acquired in 177Lu, 131I, and 225Ac imaging respectively) were simulated using the SIMIND Monte Carlo program [37] at 1100 positions that linearly varied from 0 cm to 58.44 cm.
 The detector pixel size was $0.24~{}\mathrm{cm}\times 0.24~{}\mathrm{cm}$ with $255\times 255$ pixels.
-The simulated collimator corresponded to a Siemens high energy configuration with a hexagonal shape, a hole width of 0.40.40.4~{}0.4cm and a hole length of 5.975.975.97~{}5.97cm.
+For 177Lu, the simulated collimator corresponded to a Siemens low energy high resolution (LEHR) configuration.
+While typical clinical protocols use a medium energy (ME) configuration to minimize septal penetration and scatter, the choice of collimator here is merely to demonstrate the feasibility of using a LEHR collimator with an advanced PSF model in reconstruction.
+A Siemens high energy (HE) configuration was used for 131I and 225Ac, which is standard for clinical protocol.
 The intrinsic resolution of the detector was also included in the simulation and was assumed to be 0.380.380.38~{}0.38cm at 140 keV, representative of a Symbia system with $3/8$” crystal length.
-The PSFs at all 1100 distances were used to generate $K_{\mathrm{true}}^{x,y;d}$ in Equation 19 during reconstruction by selecting the PSF closest to the actual source-detector distance.
-Twelve of these PSFs at source detector distances of 111~{}1cm and every 555~{}5cm from 555~{}5cm to 555555~{}55cm were used as data to fit the 1D-R model.
+The collimator cover and backscatter components of the detector were modeled via 0.1 cm of aluminum and 6.6 cm of pyrex respectively; inclusion of these parameters is necessary to capture further scattering of detected photons present in real systems.
+Twelve of the simulated PSFs per isotope/collimator configuration (at source-detector distances of 111~{}1cm and every 555~{}5cm from 555~{}5cm to 555555~{}55cm) were used as data to fit the 1DR-MC model.
 Fitting was performed in three steps: (i) the Gaussian parameters were optimized using the Adaptive Moment Estimation (ADAM) algorithm for $1\cdot 10^{4}$ iterations and a learning rate of $10^{-2}$, (ii) all the parameters were simultaneously optimized using ADAM for $1.5\cdot 10^{4}$ iterations and a learning rate of $10^{-3}$.
-The computational time for the 1D-R model (Equation 8) and 2D model (Equation 19) were bench-marked on CPU and GPU for by applying the operators to matrices of four different sizes: $64^{3}$, $128^{3}$, $196^{3}$, and $256^{3}$.
+The PSFs at all distances were used to build the 2D-MC models.
+The computational time for the 1DR-MC and 2D-MC models were bench-marked on CPU and GPU for by applying the 225Ac operators to matrices of four different sizes: $64^{3}$, $128^{3}$, $196^{3}$, and $256^{3}$.
 Each matrix of size $N^{3}$ was filled with random uniform numbers between 0 and 1.
@@
 2.3 SPECT Studies
-2.3.1 Phantom Study
-Application of (i) standard Gaussian (ii) 1D-R (iii) 2D PSF modeling was evaluated for GPU reconstruction of acquired 225Ac data.
+2.3.1 Monte Carlo Phantom Studies
+To demonstrate the versatility of the proposed method, application of the 2DG, 2D-MC, and 1DR-MC models were evaluated for reconstruction of MC simulated 177Lu, 131I, and 225Ac projection data.
+Each case consisted a cylindrical phantom with spheres of diameter 606060~{}60mm, 282828~{}28mm, and 222222~{}22mm, and a source to background ratio of 10:1.
+The activity concentrations in the spheres of 200200200~{}200Bq/mL (225Ac) and 111~{}1MBq/mL (177Lu and 131I) were selected based on the expected activity concentrations in patients.
+The SIMIND detector and collimator parameters were identical to the point source simulations, but data were now collected at a matrix size of $128\times 128$ pixels with $4.8~{}\mathrm{mm}\times 4.8~{}\mathrm{mm}$ resolution.
+96 projection angles at 15 s/projection were collected for 177Lu and 131I, while 32 projection angles at 45000 s/projection were collected for 225Ac; while the 177Lu and 131I configurations were representative of a typical clinical acquisition, the 225Ac configuration was such that the number acquired counts was significantly higher.
+The energy windows used for TEW scatter correction of each isotope are shown in Table 1.
+Projection data were reconstructed with OSEM for up to 20 iterations with 8 subsets (177Lu and 131I) and up to 200 iterations with 1 subset (225Ac).
+Calibration factors were used to convert images to units of MBq/mL, and were obtained via corresponding point source simulations with the same detector configuration.
+The background region used to compute contrast corresponded to four 30 mm spherical ROIs placed in the warm region.
+Table 1: Energy windows employed for all image acquisitions and reconstructions in this paper to the nearest decimal in keV.
+  • Isotope Photopeak Lower Upper
+    177Lu (MC Phantom) 187.2-228.8 169.4-187.2 228.8-252.9
+    131I (MC Phantom) 332.0-405.0 310.0-332.0 405.0-427.0
+    225Ac (MC Phantom) 396.0-484.0 374.0-396.0 484.0-506.0
+    225Ac (Physical Phantom) 396.0-484.0 352.0-396.0 484.0-572.0
+    225Ac (Patient) 396.0-484.0 358.2-395.9 484.0-506.0
+2.3.2 Physical Phantom Studies
+Application of the 2DG, 2D-MC, and 1DR-MC models were evaluated for reconstruction of acquired 225Ac physical phantom data.
 A cylindrical phantom with spheres of diameter 606060~{}60mm, 282828~{}28mm, and 222222~{}22mm with an initial sphere activity concentration of 1.37 kBq/mL was filled at a 10:1 source to background ratio.
-34 SPECT acquisitions of the phantom were taken in sequence on a Symbia T2 SPECT/CT system (Siemens Healthineers, USA) with the following settings: $128\times 128$ pixels at $4.82~{}\mathrm{mm}\times 4.82~{}\mathrm{mm}$ resolution, 96 projection angles, high energy collimators, and 60 s acquisition time per projection.
-For reconstruction of the 440440440~{}440keV emission from the 213Bi daughter, two energy windows were configured for acquisition: (i) photopeak window centered at 440 keV with a width of 20% and (ii) a lower scatter window used for DEW scatter correction centered at 374374374~{}374keV with a width of 12.5%.
-A blank scan with identical parameters was acquired to obtain the mean-stray radiation noise in each energy window.
+34 SPECT acquisitions of the phantom were taken in sequence on a Symbia T2 SPECT/CT system (Siemens Healthineers, USA) with the following settings: $128\times 128$ pixels at $4.8~{}\mathrm{mm}\times 4.8~{}\mathrm{mm}$ resolution, 96 projection angles, high energy collimators, and 60 s acquisition time per projection; although 96 projection angles were acquired, only 32 were used since this was similar to the corresponding 225Ac patient acquisition in the next section.
+The energy windows selected for TEW additive correction are shown in Table 1.
+A blank scan with identical parameters was acquired to obtain the mean stray radiation noise in each energy window.
+Images were reconstructed with MLEM for up to 200 iterations.
 Two noise levels of the data were considered:
@@
     One of the 34 scans was used for data reconstruction.
-The number of acquired counts in this scenario corresponds to an expected clinical 225Ac scan of a patient undergoing 225Ac-PSMA-617 therapy with injection of 888~{}8MBq [23] and is scanned for 2.5 min per projection at a time point between 0-72 hr post injection.
+The number of acquired counts in this scenario is approximately representative of a clinical 225Ac-PSMA-617 scan, where the patient is injected with 888~{}8MBq of activity [23] and scanned for 2.5 min per projection at some time point between 0 hr and 72 hr post injection.
   2. 2.
     The counts from all 34 scans were summed together and used for image reconstruction.
-In this case, no scatter blurring was used in Equation 21 since the noise level was low.
 Because the detectors had an approximately equal radial path around the phantom for each scan, this corresponds to a single scan with a high count rate.
-Although 96 projection angles were acquired, only 32 were used since this was similar to the corresponding 225Ac patient acquisition in the next section.
-All images were reconstructed with MLEM for up to 100 iterations.
+Spherical masks of each hot sphere were drawn using bright edges on the CT image (visible in Figure 5) so that the counts could be computed for each iteration number.
+Three additional spherical masks of diameter 50 mm were drawn in the background region in order to compute contrast.
 2.3.3 Patient Study
@@
 SPECT data were acquired on a Discovery 670 Pro SPECT/CT (GE Healthcare, USA) with a high energy general purpose collimator.
-30 projections (15 per head) were acquired for 150 s using an identical photopeak window to the phantom acquisition, and 358.15 keV - 395.85 keV lower window.
-Since this GE scanner has a different CDR than the Siemens scanner used in prior examples, PSF data were regenerated in SIMIND using a high energy general purpose GE collimator and another 1D-R model was fit as before.
+30 projections (15 per head) were acquired for 150 s with energy windows given by Table 1.
+Since this GE scanner has a different CDR than the Siemens scanner used in prior examples, PSF data were regenerated in SIMIND using a high energy general purpose GE collimator and another 1DR-MC model was fit as before.
 In GE scanners, the hexagonal bores in the collimator are rotated by $90^{\circ}$ compared to the Siemens collimator, so the angles in $\Theta_{t}$ were adjusted to compensate for this.
-The acquired data were reconstructed using (i) standard Gaussian (ii) 2D and (iii) 1D-R PSF modeling using MLEM for up to 100 iterations.
-Reconstructed images were filtered using a Gaussian filter with a full width half max (FWHM) of 333~{}3cm.
-Two high uptake bone lesions were segmented on the 100th iteration reconstruction using 3D Slicer [39].
+The acquired data were reconstructed using 2DG, 2D-MC, and 1DR-MC PSF modeling and MLEM for up to 200 iterations.
 Three lesion ROIs were segmented by a physician on a pre-therapeutical PET image using the PET Edge+ tool of MIM v7.2.1 (MIM Software Inc., USA).
 A spherical ROI of diameter 101010~{}10cm was placed in a low uptake region in the center of the patient so that the contrast, defined as the mean uptake ratio between the lesion ROIs and background ROI, could be obtained.
-The mean number of counts and contrast for each lesion ROI were then evaluated for each iteration of MLEM; before the statistics were computed, the 333~{}3cm FWHM Gaussian filter was applied to the image.
+The mean number of counts and contrast for each lesion ROI were then evaluated for each iteration of MLEM; in this case, before the statistics were computed for each iteration, a 333~{}3cm FWHM Gaussian filter was applied to the image.
 3 Results
-Figure 2 shows the Monte Carlo simulated PSF data compared to the fit obtained via Equation 8.
-The curve fit is a reasonable approximation to the PSF at most distances, but is unable to capture features at larger distances, such as the intensity of the tails at 50 cm.
-The model also consistently under predicts the intensity at the center of the PSF, and this effect is more prevalent at large distances.
-Figure 2: Comparison between 225Ac Monte Carlo simulated PSF data and the corresponding fit from the 1D-R model at different source-detector distances.
-From top to bottom, (i) 2D profiles of Monte Carlo PSF data, (ii) 2D profiles of 1D-R model, (iii) central vertical profile of Monte Carlo and fitted PSFs, (iv) central horizontal profile of Monte Carlo and fitted PSFs; the Monte Carlo data is shown in black scatter points, while the fitted data is shown as a solid red line.
-Timing benchmarks for PSF modeling using the 1D-R, 1D-R (FFT), 2D, and 2D (FFT) are shown in Figure 3.
-CPU implementation yields no benefits with the proposed 1D-R model, and performs fastest with FFT-based 2D convolutions and slowest using regular 2D convolutions.
-The GPU implementation is faster than the CPU implementation for all methods, and yields computational benefits when the proposed 1D-R model is used.
-With a matrix size of $128^{3}$, use of the 1D-R method is over three times faster than 2D (FFT) PSF modeling.
-For small matrix sizes ($64^{3}$) there is no computational speed-up using the 1D-R method, and for large matrix sizes ($256^{3}$) the relative time difference between the 1D-R and 2D (FFT) begins to decrease, approaching only a two times speed advantage.
-Figure 3: CPU and GPU required run time of the 1D-R and 2D PSF models for different matrix sizes.
+Figure 2 compares the Monte Carlo simulated 2D-MC PSF profiles to the 1DR-MC fit obtained via Equation 8 for 177Lu, 131I, and 225Ac.
+The 1D-R model reasonable approximates all PSFs, but (i) does not always accurately capture the septal scatter component as seen in the central vertical profiles and (ii) does not capture some additional faint septal penetration tails of 177Lu; this is due to a limitation of the model for only including three tails at specific angles.
+Figure 2: Comparison between Monte Carlo simulated PSF data and the corresponding fit from the 1DR-MC model at different source-detector distances.
+From top to bottom, (i) 2D profiles of Monte Carlo PSF data, (ii) 2D profiles of 1DR-MC model, (iii) central vertical profile of Monte Carlo and fitted PSFs, (iv) central horizontal profile of Monte Carlo and fitted PSFs; the Monte Carlo data is shown in black scatter points, while the fitted data is shown as a solid red line.
+Timing benchmarks for PSF modeling using the 1DR-MC, 1DR-MC (FFT), 2D-MC, and 2D-MC (FFT) are shown in Figure 3.
+CPU implementation yields no benefits with the proposed 1DR-MC model, and performs fastest with FFT-based 2D convolutions and slowest using regular 2D convolutions.
+The GPU implementation is faster than the CPU implementation for all methods, and yields computational benefits when the proposed 1DR-MC model is used.
+With a matrix size of $128^{3}$, use of the 1DR-MC method is over three times faster than 2D-MC (FFT) PSF modeling.
+For small matrix sizes ($64^{3}$) there is no computational speed-up using the 1DR-MC method, and for large matrix sizes ($256^{3}$) the relative time difference between the 1DR-MC and 2D-MC (FFT) begins to decrease, approaching only a two times speed advantage.
+Figure 3: CPU and GPU required run time of the 1DR-MC and 2D PSF models for different matrix sizes.
 Each implementation is evaluated using standard and FFT based convolution.
 The time shown on top of each bar in units of ms.
-PyTomography reconstructed 225Ac phantom images are shown in Figure 4.
-The time required for 100 iterations of MLEM for the low count data was 110110110~{}110s (Gaussian PSF), 146146146~{}146s (1D-R PSF) and 377377377~{}377s (2D PSF).
-Reconstructions using the 1D-R and 2D methods are almost indistinguishable qualitatively.
-All 34 low count acquisitions were then reconstructed and the 1D-R and 2D methods were quantitatively compared; the differences between the 2D and 1D-R were $(1.25\pm 1.25)\%$, $(0.60\pm 0.69)\%$, and $(1.90\pm 0.18)\%$ from smallest to largest sphere respectively.
-For comparison, the variability in counts between separate noise realizations for the 2D model were $54.4\%$, $33.0\%$, and $9.8\%$ from smallest to largest sphere, respectively.
-Figure 4: Axial slices from reconstruction of low-count (top row) and high-count (bottom row) 225Ac phantom data acquired on a Symbia T2 SPECT system.
+Reconstructed MC simulations of the 177Lu, 131I, and 225Ac spherical phantoms are shown in Figure 4, while corresponding RC and contrast statistics (for the last iteration of reconstruction) are shown in Table 2.
+The 1DR-MC and 2D-MC models yield similar RCs in all cases, and are larger than those obtained using the 2DG model.
+Figure 4: Reconstructions of the MC simulations (left) and corresponding RCs for each of the three spheres at each iteration (right).
+The displayed images correspond to reconstructions after 40 iterations (177Lu), 10 iterations (131I) and 100 iterations (225Ac)
+Table 2: Recovery coefficients (or counts) and contrast for various regions of interest in all reconstruction examples using (i) traditional 2D Gaussian PSF modeling (2DG), (ii) 2D kernels (2D-MC), and (iii) the proposed 1DR-MC model that aims to to be consistent with the 2D-MC model while reducing reconstruction times.
+The low count 225Ac physical phantom corresponds to the first of the 34 acquisitions.
+  • 177Lu MC Phantom
+    Contrast RC [%]
+    Model 60 mm 28 mm 22 mm 60 mm 28 mm 22 mm
+    2DG 2.68 2.00 1.73 28.4 21.2 18.3
+    2D-MC 7.85 5.72 4.23 95.9 69.9 51.7
+    1DR-MC 8.19 6.01 4.40 94.6 69.4 50.8
+    131I MC Phantom
+    Contrast RC [%]
+    Model 60 mm 28 mm 22 mm 60 mm 28 mm 22 mm
+    2DG 8.08 6.94 6.22 75.4 64.7 58.0
+    2D-MC 10.41 9.60 8.74 94.1 86.7 79.0
+    1DR-MC 10.31 9.49 8.64 94.0 86.5 78.8
+    225Ac MC Phantom
+    Contrast RC [%]
+    Model 60 mm 28 mm 22 mm 60 mm 28 mm 22 mm
+    2DG 4.55 3.73 3.63 45.5 37.3 36.2
+    2D-MC 9.11 8.42 7.05 98.8 91.2 76.5
+    1DR-MC 9.03 8.36 6.99 98.7 91.4 76.5
+    225Ac Physical Phantom (Low Count)
+    Contrast Counts
+    Model 60 mm 28 mm 22 mm 60 mm 28 mm 22 mm
+    2DG 4.01 3.68 3.32 0.0342 0.0314 0.0283
+    2D-MC 4.85 4.80 4.36 0.0640 0.0634 0.0575
+    1DR-MC 4.88 4.80 4.36 0.0643 0.0632 0.0574
+    225Ac Physical Phantom (High Count)
+    Contrast Counts
+    Model 60 mm 28 mm 22 mm 60 mm 28 mm 22 mm
+    2DG 4.68 4.31 3.70 1.24 1.14 0.98
+    2D-MC 8.15 7.83 5.98 2.52 2.42 1.85
+    1DR-MC 8.05 7.69 5.92 2.56 2.45 1.88
+    225Ac Patient
+    Contrast Counts
+    Model Les 1 Les 2 Les 3 Les 1 Les 2 Les 3
+    2DG 2.32 2.28 2.80 0.0405 0.0267 0.0660
+    2D-MC 3.72 3.50 4.86 0.0405 0.0273 0.0653
+    1DR-MC 3.74 3.50 4.92 0.0246 0.0254 0.0355
+Reconstructions of physical 225Ac phantom images are shown in Figure 5, and quantitative statistics for (i) the first low count acquisition and (ii) the high count acquisition after 200 iterations are shown in Table 2.
+Images generated from the 1DR-MC and 2D-MC methods are almost indistinguishable qualitatively.
+The quality metrics in Table 2 demonstrate similar counts and contrast for the 2D-MC and 1DR-MC models, and significantly lower counts and contrast for the 2DG model.
+The time required for 200 iterations of MLEM for the low count data was 777777~{}77s (2DG PSF), 299299299~{}299s (1DR-MC PSF) and 748748748~{}748s (2D-MC PSF).
+All 34 low count acquisitions were also reconstructed and the 2DG, 2D-MC, and 1DR-MC methods were quantitatively compared; the differences in total counts between the 2D-MC and 1DR-MC were $(-0.01\pm 1.20)\%$, $(0.67\pm 0.23)\%$, and $(-0.75\pm 1.25)\%$ for smallest to largest sphere respectively.
+As a demonstration of how minor these differences are, the variability of counts between the 34 acquisitions for the 2D-MC model were $61.3\%$, $38.5\%$, and $11.1\%$ from smallest to largest sphere, respectively.
+Figure 5: Reconstruction of physical 225Ac phantom using various PSF models.
 Reconstructed SPECT and CT images are superimposed; the boundaries of each sphere can be seen on the CT.
-Images were reconstructed with the reference 2D, Gaussian, and proposed 1D-R PSF modeling using the acquired Monte Carlo data.
-The displayed low count images were post-smoothed using a 3D Gaussian function with a 3 cm FWHM.
-The plots on the right show the average counts per voxel in the small (sml) medium (med) and large (lrg) spheres as a function of iteration number; the 2D is shown in black, the Gaussian is shown in blue, and the 1D-R is shown in red.
-PyTomography reconstructed 225Ac-PSMA-617 patient images are shown in Figure 5.
-The time required for 100 iterations of MLEM was $38.5$ s (Gaussian PSF), 165.6165.6165.6~{}165.6s (1D-R PSF) and 351.5351.5351.5~{}351.5s (2D PSF).
-Qualitatively, the Gaussian PSF yields less counts in the uptake ROIs, and also greater background counts, for example, in the vertebrae.
-The count differences between the 1D-R and 2D methods after 100 iterations were 1.271.271.27~{}1.27(lesion 1), 0.480.480.48~{}0.48(lesion 2) and $2.03~{}\%$ (lesion 3).
-These differences are comparable to those observed in the phantom study, and were significantly less than the differences between the Gaussian and 2D methods of $-42.3~{}\%$ (lesion 1), $-29.1~{}\%$ (lesion 2) and $-44.3~{}\%$ (lesion 3).
-The differences in lesion-background contrast between the 1D-R and 2D methods after 100 iterations were $-0.18~{}\%$ (lesion 1), $-0.96~{}\%$ (lesion 2) and $0.56~{}\%$ (lesion 3).
-These were smaller in magnitude than the contrast differences between the Gaussian and 2D methods of $-51.3\%$ (lesion 1) and $-23.3~{}\%$ (lesion 2) and $-56.7~{}\%$ (lesion 3).
-Based on the line plots in Figure 5, the Gaussian PSF yields convergence with less iterations; this observation is similar to how an absence CDR modeling leads to faster convergence in traditional low and medium energy SPECT.
-In general, the smaller the spatial extent of the PSF, the faster convergence is attained in MLEM.
-Figure 5: Coronal slices of 68Ga-PSMA-617 (top row, left) and 225Ac-PSMA-617 patient data (top row, right), counts in bone lesion ROIs as a function of iteration number (middle row) and signal to background ratios in the bone lesion ROIs (bottom row).
-The Background ROI was a 10 cm sphere placed in the center of the patient with little activity.
-The 68Ga-PSMA-617 image was reconstructed using QClear on the vendor’s software.
-225Ac images were reconstructed in PyTomography using (i) reference 2D, (ii) Gaussian, and (iii) the proposed 1D-R model with 100 iterations of MLEM.
-The displayed reconstructed SPECT images were post filtered using a 3D 3 cm FWHM Gaussian.
+The displayed images correspond to reconstruction after 100 iterations; the low count images were additionally post-smoothed using a 3D Gaussian function with a 3 cm FWHM.
+The plots on the right show the average counts per voxel in the three spheres.
+PyTomography reconstructed 225Ac-PSMA-617 patient images are shown in Figure 6, along with a reference 68Ga-PSMA-617 PET scan acquired on a GE Discovery IQ scanner and reconstructed using Q.Clear ($\beta=450$).
+The time required for 200 iterations of MLEM was $75.1$ s (2DG), 272.0272.0272.0~{}272.0s (1DR-MC) and 704.0704.0704.0~{}704.0s (2D-MC).
+Qualitatively, images reconstructed using the Gaussian PSF can be observed to have less counts in the lesion ROIs.
+Quantitative results are shown as line plots in Figure 6, and quality metrics for iteration 200 are shown in Table 2.
+These results demonstrate (i) indistinguishable images and nearly identical image metrics between the 1DR-MC and 2D-MC models and (ii) higher contrast and count uptake for 2D-MC and 1DR-MC models relative to the 2DG model for all three lesions.
+Figure 6: Coronal slice of a 68Ga-PSMA-617 PET image (top row, left) and associated 225Ac-PSMA-617 SPECT data (top row, right), counts in bone lesion ROIs as a function of iteration number (middle row) and signal to background ratios in the bone lesion ROIs (bottom row).
+225Ac images were reconstructed in PyTomography using the 2DG, 2D-MC, and 1DR-MC PSF models.
 For computation of the data in the line plots, the image at each iteration was post-filtered using a 3 cm FWHM 3D Gaussian filter before the statistic was computed.
@@
 4 Discussion
-In this work, a 1D-R PSF model for high energy SPECT reconstruction was proposed, and demonstrated to be both efficacious and fast for use in high-energy reconstruction of 225Ac data.
-The model was created using publicly shared SPECTPSFToolbox component of the PyTomography project; t is hoped that the shared toolbox and corresponding tutorials will permit the nuclear medicine community to develop custom PSF models for new isotope/collimator configurations as research in novel RPTs and imaging techniques continues to grow.
-As shown in Figure 3, the proposed 1D-R only yields computational benefits when image reconstruction is performed on GPU.
-If reconstruction is performed on CPU, then PSF modeling should be implemented with FFT-based 2D convolutions.
-FFT-based 2D convolutions are still, however, three times faster on GPU than on CPU.
-Furthermore, if a GPU is used then the 1D-R model can be used for further computational advantages.
-The fastest implementation of PSF modeling on GPU (proposed 1D-R) is approximately one order of magnitude faster than the fastest implementation of PSF modeling on CPU (FFT-based 2D convolutions).
-While the proposed 1D-R model is not able to perfectly capture the features of the PSF at all radial distances in Figure 2, it yields near identical reconstructions in Figures 4 and 5, especially when compared to reconstruction using the Gaussian kernel.
-In the low count scenario of the phantom study, which approximately represents a clinical count rate, the percent difference in the three spheres between the 1D-R and 2D model were small compared to the variability between separate acquisitions.
-In the patient example, the variability of counts between the 2D and 1D-R models in the bone lesions was similar to the spheres in the phantom: this suggests that any differences in predicted counts between the 2D and 1D-R models in the patient example is also negligible compared to differences that would be observed between separate acquisitions.
-While the 1D-R model and 2D model yielded similar contrast, the Gaussian model had significantly reduced contrast with differences of $-51.3~{}\%$, $-23.3~{}\%$, and $-56.7~{}\%$ compared to the 2D model in the bone lesions.
-While scaling the magnitude of the image reconstructed using the Gaussian PSF could artificially increase the number of counts in each bone ROI, it would have no impact on these substantial differences in contrast.
-In the phantom example, the radial path of the detector approached 7.27.27.2~{}7.2cm at its minimum and 27.227.227.2~{}27.2cm at its maximum.
-Since the radius of the cylinder was 101010~{}10cm, this suggests that the PSF model was able to reasonably function between $0-37~{}$cm.
-Future work may seek to test the efficacy of the proposed 1D-R model for larger radial distances, as this may be required for larger patients.
-While the functional form of Equations 9, 12, 15 were selected because they produced a reasonable approximation of the PSF, it may be the case that they can be improved upon.
+This work proposes a 1DR-MC PSF model for high-energy SPECT reconstruction that demonstrates consistency with 2D-MC models, while also increasing computational efficiency for GPU-based reconstruction of 177Lu, 131I, and 225Ac at typical SPECT matrix sizes.
+The model was created using the publicly shared SPECTPSFToolbox component of the PyTomography project; it is hoped that the shared toolbox and corresponding tutorials will permit the nuclear medicine community to develop custom PSF models for new isotope/collimator configurations as research in novel RPTs and imaging techniques continues to grow.
+As shown in Figure 3, the proposed 1DR-MC only yields computational benefits when image reconstruction is performed on GPU.
+If reconstruction is performed on CPU, then PSF modeling is faster when implemented with FFT-based 2D convolutions.
+As discussed in Section 2.1.2, the kernel sizes were fixed to $N-1\times N-1$ for all $d$.
+For small $d$, where the PSF is also smaller, a kernel of this size might be excessive and could unnecessarily increase the computational time required.
+Future research exploring the reduction of computational time might experiment with using a source-detector distance dependent kernel size $N(d)$ for that matches the size of the PSF; this would further reduce the computational time required.
+While the proposed 1DR-MC model is not able to perfectly capture all the PSF features in Figure 2, it yields qualitatively similar reconstructions of the MC data in Figure 4.
+Furthermore, as shown in Table 2, the RCs and contrast of the 1DR-MC and 2D-MC models are similar and always significantly greater than the 2DG model.
+This demonstrates the superior quantitative performance of the MC-based models, and justification for using the 1DR-MC over the 2D-MC method due to faster reconstruction times.
+For the physical 225Ac phantom and patient use cases shown in Figures 5 and 6, the 2D-MC and 1DR-MC models yield similar count quantification, especially when compared to the 2DG model.
+The quality metrics for 200 iterative updates shown in Table 2 further highlight this consistency.
+In the physical phantom high count case, use of the 1DR-MC method compared to the 2DG method increased the mean counts the spheres by factors of 2.06, 2.14, and 1.92 for the largest to smallest sphere respectively; this is somewhat similar to the MC study where the RCs increased by 2.17, 2.45, and 2.11 for spheres of the same size.
+While calibration factors were not available for the physical data, inference from the MC results suggest that the RCs of the physical data were significantly improved.
+In the low count scenario of the physical phantom study, which approximately represents a clinical count rate, the percent difference in the three spheres between the 2D-MC and 1DR-MC models were small compared to the variability between the thirty-four separate acquisitions.
+This highlights that any systematic errors between use of the 2D-MC and 1DR-MC model are negligible compared to the precision of the imaging protocol, and thus that the 1DR-MC model may be suitable for standard 225Ac image reconstruction.
+In the patient example, it was similarly found that (i) the differences between the 2D-MC and 1DR-MC models in the bone lesions were small and (ii) the 2DG model had significantly reduced contrast relative to the 2D-MC and 1DR-MC models.
+The functional form of Equations 9, 12, and 15 were selected because they produced a reasonable approximation of the PSF, but they can be further developed.
 Use of rotation operations with the component $K_{B}^{(x;d;b)}K_{B}^{(y;d;b)}$ may permit more radial symmetry in the SSRF.
-Substitution of $K_{G}^{(x;d;b)}K_{G}^{(y;d;b)}$ with a small 2D kernel that represents the true aperture function of the GRF may remedy the under prediction in the center of the PSFs at large distances.
-Close inspection of the SIMIND data at large distances in Figure 2 reveals additional dim tails; these could be modeled by including another independent tail component in the model.
+Substitution of $K_{G}^{(x;d;b)}K_{G}^{(y;d;b)}$ with a small 2D kernel that represents the true aperture function of the GRF may also improve the accuracy of the PSF model.
+The 177Lu PSF data in Figure 2 reveals additional dim tails; these could be modeled by including another independent tail component in the model.
 A trade-off for adding additional features to the model, however, is an increase in computation time.
@@
 Users can independently evaluate the trade-off between adding model features, impact on reconstructed images, and increase in computation time.
-As discussed in Section 2.1.2, the kernel sizes were fixed to $N-1\times N-1$ for all $d$.
-For small $d$, where the PSF is also smaller, a kernel of this size might be excessive and could unnecessarily increase the computational time required.
-Future research exploring the reduction of computational time might experiment with using a source-detector distance dependent kernel size $N(d)$ for that matches the size of the PSF; this would further reduce the computational time required.
-Since reconstruction using the proposed 1D-R model more than halves reconstruction times compared to the 2D model, it may be preferable for reconstruction of 225Ac data.
-Both the 1D-R and 2D models, however, are only applicable when Equation 5 is used for SPECT system matrix modeling; this equation relies on the assumption that the attenuation probabilities vary little across the PSF.
-Due to the large spatial extent of the PSFs shown here, this assumption may be invalid, and Equation 4 may instead be required for accurate reconstruction of 225Ac data.
+Since GPU-accelerated reconstruction using the proposed 1DR-MC model reduces reconstruction times by more than a factor of two compared to the 2D-MC model, it may be preferable for reconstruction of 225Ac data for typical SPECT matrix sizes.
+Both the 1DR-MC and 2D-MC models, however, are only applicable when Equation 5 is used for SPECT system matrix modeling; this equation relies on the assumption that the attenuation probabilities vary little across the PSF.
+Due to the large spatial extent of the PSFs shown here, this assumption may be invalid, and Equation 4 may instead be required for accurate reconstruction.
 This assumption largely depends on the density of the object being scanned.
@@
 Meanwhile, the study of different density configurations and the applicability of Equation 5 remains to be investigated in future work.
-In conclusion, a fast and accurate implementation of high energy CDR modeling in SPECT imaging that uses 1D convolutions and rotations was developed, and tested on 225Ac reconstructions.
-The technique was implemented using the open source reconstruction library PyTomography, and was shown to speed up reconstruction times by more than a factor of two compared to conventional, 2D convolution based methods.
-Furthermore, the 1D-R method yielded near identical results to the 2D method, and the small differences were insignificant compared to the differences between reconstructions of separate noise realizations.
+5 Conclusion
+In conclusion, a computationally efficient implementation of high energy CDR modeling in SPECT imaging was developed and tested on 177Lu, 131I, and 225Ac reconstructions.
+The technique was implemented using the open source reconstruction library PyTomography, and was shown to speed up reconstruction times by more than a factor of two compared to conventional, 2D convolution based methods ($272.0$ s vs.
+$704.0$ s for patient data).
+Furthermore, the proposed 1DR-MC method yielded near identical results to the reference 2D-MC approach, and the small differences for 225Ac were insignificant compared to the differences between reconstructions of separate noise realizations.
 The SPECTPSFToolbox python library was developed and publicly shared to permit others in the nuclear medicine community to develop custom isotope/collimator PSF models for use in the open-source reconstruction library PyTomography.